### v4

# Heterochromatic Geometric Transversals of Convex sets

Abstract
An infinite sequence of sets $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence from an infinite sequence of families $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$, if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. In this paper, we have proved that if for each $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is a family of nicely shaped convex sets in $\mathbb{R}^{d}$ such that each heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ contains at least $k+2$ sets that can be pierced by a single $k$-flat ($k$-dimensional affine space) then all but finitely many families in $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ can be pierced by finitely many $k$-flats. This result can be considered as a countably colorful generalization of the $(\aleph_{0},k+2)$-theorem proved by Keller and Perles (Symposium on Computational Geometry 2022). We have also established the tightness of our result by proving a number of no-go theorems.
1 Introduction
The following theorem due to Helly [Hel23], proved more than 100 years ago, has been one of the fundamental results in discrete and convex geometry with numerous extensions and applications.
Theorem 1.1 (Helly [Hel23]).
Let $\mathcal{F}$ be an infinite family of compact convex sets in $\mathbb{R}^{d}$ such that the intersection of every $d+1$ sets from $\mathcal{F}$ is non-empty then the intersection of all the sets in $\mathcal{F}$ is non-empty.
For a comprehensive introduction to Helly’s Theorem and related topics see the recent surveys by Amenta, Loera, and Soberón [ADLS17], and Bárány and Kalai [BK22]. Unless otherwise stated explicitly, whenever we talk about a family of convex sets we won’t assume anything about the cardinality of the family, that is, a family of convex sets can be a finite collection or an infinite collection.
A $k$-flat $H$ is said to hit or pierce a set $F\subseteq\mathbb{R}^{d}$ if $F\cap H\neq\emptyset$. A set $\mathcal{T}$ of $k$-flats is called a $k$-transversal of a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ if for every $F\in\mathcal{F}$ there exists a $k$-flat in $\mathcal{T}$ that pierces $F$. Additionally, if $k=0$ or $k=1$ or $k=d-1$ then $\mathcal{T}$ is called point transversal, line transversal and hyperplane transversal of $\mathcal{F}$ respectively. Also, we say a family $\mathcal{F}$ has a $k$-transversal of size $m$ (or, at most $m$) if there exists a $k$-transversal $\mathcal{T}$ of $\mathcal{F}$ and $|\mathcal{T}|=m$ (or, $|\mathcal{T}|\leq m$). Generalizations of Helly’s Theorem to general $k$-transversals, for $11\leq k\leq d-$ 1, has been an active area of research, see the survey by Holmsen and Wenger [HW17]. Santaló [San40] showed that there can be no analog of Helly Theorem for arbitrary families of convex sets and line transversals. Because of Santaló’s impossibility result, the follow-up works on $k$-transversals are for families of convex sets with additional properties, for example, families of unit balls, etc. Hadwiger [Had56] proved the following Helly-type result for line transversal.
Theorem 1.2 (Hadwiger [Had56]).
Let ${C}$ be a compact convex set in $\mathbb{R}^{d}$ with a non-empty interior, and $\mathcal{F}$ be a countably infinite collection of disjoint convex bodies congruent to $\mathcal{C}$. If every subset of $\mathcal{F}$ of size $d+1$ has a line transversal of size one then the whole family $\mathcal{F}$ has a line transversal of size one.
Danzer, Grünbaum and Klee [DGK63] showed that the "congruent" assumption in Theorem 1.2 can be weakened as long as the sets are of bounded diameter, that is, there exists $R>0$ such that the diameter of all the convex sets in the family is at most $R$.
Theorem 1.3 (Danzer, Grünbaum, and Klee [DGK63]).
Suppose $R>0$ and $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ with diameters at most $R$ and $\bigcup_{C\in\mathcal{F}}C$ is unbounded. If every subset of $\mathcal{F}$ of size $d+1$ has a line transversal of size one then the whole family $\mathcal{F}$ has a line transversal of size one.
A family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is said to satisfy $(p,q)$-property with respect to $k$-flats, for $10\leq k\leq d-$ 1, if every subset of $\mathcal{F}$ of size $p$ contains $q$ sets that can be pierced by a single $k$-flat. Hadwiger and Debrunner [HD57] were the first to introduce $(p,q)$-property, and they proved that for all $p\geq q\geq d+1$ satisfying $p(d-1)<(q-1)d$, if $\mathcal{F}$ is a family of compact convex sets in $\mathbb{R}^{d}$ satisfying $(p,q)$-property with respect to points then $\mathcal{F}$ has a point transversal of size at most $p-q+1$. Hadwiger and Debrunner [HD57] asked the following question that remained open for $35$-years.
Question 1.4 (Hadwiger and Debrunner $(p,q)$-problem [HD57]).
For all $p\geq q\geq d+1$, does there exists a constant $c(p,q,d)>0$ such that if a family $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{d}$ satisfy $(p,q)$-property with respect points then $\mathcal{F}$ has a point transversal of size at most $c(p,q,d)$?
Alon and Kleitman [AK92] proved the following fundamental theorem and since then this area has been one of the most active areas of research in Discrete Geometry.
Theorem 1.5 (Alon-Kleitman [AK92]: $(p,q)$-theorem).
For any three natural numbers $p\geq q\geq d+1$, $\exists c=c(p,q,d)$ such that if $\mathcal{F}$ is a collection of compact convex sets satisfying the $(p,q)$-property in $\mathbb{R}^{d}$ then there exists a point transversal for $\mathcal{F}$ of size at most $c$.
Alon and Kalai [AK95] later proved the $(p,q)$-theorem for hyperplane transversal.
Theorem 1.6 (Alon-Kalai [AK95]: $(p,q)$-theorem for hyperplane transversal).
For any three natural numbers $p\geq q\geq d+1$, $\exists c^{\prime}=c^{\prime}(p,q,d)$ such that if $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ satisfying the $(p,q)$-property with respect to hyperplanes then there exists a hyperplane transversal for $\mathcal{F}$ of size at most $c^{\prime}$.
Unlike the case of hyperplane transversal Alon, Kalai, Matoušek, and Meshulam [AKMM02] gave an explicit construction showing the impossibility of getting a $(p,q)$-theorem, like Theorem 1.6, for $k$-transversal when $10<k<d-$ 1.
Keller and Perles introduced an interesting variation of the $(p,q)$-property. A family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is said to satisfy $(\aleph_{0},q(k))$-property with respect to $k$-flats if among every infinite sequence of sets from $\mathcal{F}$ there exist $q(k)$ sets such that these $q(k)$ sets can be pierced by a single $k$-flat. Keller and Perles [KP22] proved the following important theorem for families of $(r,R)$-fat convex sets satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats. Note that a convex set $C$ is $(r,R)$-fat if there exist two closed balls $B(c_{1},r)$ and $B(c_{2},R)$ with $B(c_{1},r)\subseteq C\subseteq B(c_{2},R).$
Theorem 1.7 (Keller and Perles [KP22]: $(\aleph_{0},k+2)$-theorem for $k$-transversal).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ is an infinite family of $(r,R)$-fat convex sets, with $0<r<R<\infty$, satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats, then the whole family $\mathcal{F}$ has a finite size $k$-transversal.
Observe that for $k$-transversal $q(k)=k+2$ is optimal since any $k+1$ nonempty subset of $\mathbb{R}^{d}$ can be pierced by a single $k$-flat. Keller and Perles proved a similar result for point transversal of families of closed balls (with no restriction on radii). They also gave an explicit example showing that the compactness assumption in Theorem 1.7 cannot be weakened. Additionally, Keller and Perles observed that, unlike $(p,q)$-property, for all $10\leq k\leq d-$ 1 there does not exist a constant $c(d,k)>0$ such that for all families of unit balls in $\mathbb{R}^{d}$ satisfying $(\aleph_{0},K+2)$-property with respect to $k$-flats then these families have a $k$-transversal of size at most $c(d,k)$.
Given $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ families of convex sets in $\mathbb{R}^{d}$, we call a collection $B_{1},\dots,B_{p}$ of $p$ convex sets colorful with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ if for all $i\in\left\{1,\dots,p\right\}$ we have $B_{i}\in\mathcal{F}_{i}$. We say $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ satisfies colorful $(p,q)$-property with respect to $k$-flats if every colorful collection of $p$ sets with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ contains $q$ sets that can be pierced by a single $k$-flat. Lovász observed the following interesting variant of Helly’s theorem, called the Colorful Helly’s theorem, in 1974, though the proof first appeared in a paper on Carathéodory’s theorem by Bárány [B8́2].
Theorem 1.8 (Bárány [B8́2]: Colorful Helly’s theorem).
Suppose $\mathcal{F}_{1},\dots,\mathcal{F}_{d+1}$ are families of convex sets in $\mathbb{R}^{d}$ satisfying the colorful $(d+1,d+1)$-property with respect to points. Then $\exists\,i\in\left\{1,\dots,d+1\right\}$ and a point in $\mathbb{R}^{d}$ that pierces all the sets in $\mathcal{F}_{i}$.
Bárány, Fodor, Montejano, Oliveros, and Pór [${}^{+}$14] proved a colorful $(p,q)$-theorem for families of convex sets in $\mathbb{R}^{d}$.
Theorem 1.9 (Bárány-Fodor-Montejano-Oliveros-Pór [${}^{+}$14]: Colorful $(p,q)$-theorem).
For any three natural numbers $p\geq q\geq d+1$, $\exists M=M(p,q,d)$ such that the following holds: if $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ be finite families of convex sets in $\mathbb{R}^{d}$ satisfying colorful $(p,q)$-property with respect to points then there exists at least $q-d$ indices $i\in\left\{1,\dots,p\right\}$ such that $\mathcal{F}_{i}$ has a point transversal of size at most $M$.
Our results
In contrast with Theorem 1.9, we shall show that no colorful $(\aleph_{0},q)$-theorem exists for $k$-flats and families of unit balls, except for the case where $k=0$. However, by expanding the definition of ‘colorful’, we can obtain a weak ‘colorful’$(\aleph_{0},k+2)$-theorem for $d>k>0$ involving families of unit balls and $k$-flats. For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of compact convex sets in $\mathbb{R}^{d}$. An infinite sequence of compact convex sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. A heterochromatic sequence is said to be a strictly heterochromatic sequence if each family has a representative in the sequence, that is, a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a strictly heterochromatic sequence if for each $n\in\mathbb{N},\;B_{n}\in\mathcal{F}_{n}$.
A sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets in $\mathbb{R}^{d}$ is said to be $k$-dependent if there exists a $k$-flat $K$ and $k+2$ distinct integers $i_{1},\dots,i_{k+2}$ such that for all $j\in[k+2]$ we have $K\cap B_{i_{j}{}}\neq\emptyset.$ If a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets is not $k$-dependent then we will call it $k$-independent. A collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of families of convex sets in $\mathbb{R}^{d}$ is said to satisfy heterochromatic $(\aleph_{0},k+2)$-property if every heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent. Similarly, we will say that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies weak heterochromatic $(\aleph_{0},k+2)$-property if every strictly heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
In this paper, we give a number of countably colorful $(\aleph_{0},k+2)$-theorems for $k$-transversals.
Theorem 1.10 (Countably colorful $(\aleph_{0},\cdot)$-theorem).
Let $d\in\mathbb{N}$, and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of $(r,R)$-fat closed convex sets in $\mathbb{R}^{d}$.
  • (i)
    For a non-negative integer $k$ with $10\leq k\leq d-$ 1 if $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the heterochromatic $(\aleph_{0},k+2)$-property, then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many $k$-flats.
  • (ii)
    If $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property, then there exist at least two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ with $i\neq j$ and each of them can be pierced by finitely many points.
The above theorem is a colorful generalization of Theorem 1.7 by Keller and Perles [KP22].
The following proposition shows that Theorem 1.10 $(ii)$ is tight for families satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Proposition 1.11.
For each $n\in\mathbb{N}$, there exists a family $\mathcal{F}_{n}$ of units balls in $\mathbb{R}^{2}$ such that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the following
  • (i)
    $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property, and
  • (ii)
    exactly two distinct $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ are pierceable by finite number of points.
Given Theorem 1.10 $(ii)$ it is natural to ask the following question:
Let $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ be an infinite sequence of families of $(R,r)$-fat closed convex sets satisfying weak heterochromatic $(\aleph_{0},k+2)$-property for $d>k>0$. Does there exist a family $\mathcal{F}_{m}$ from the sequence $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ that has a finite size $k$-transversal?
We show that the above statement is not true for all non-negative $k$ with $11\leq k\leq d-$ 1 even if we restrict ourselves to the case of unit balls.
Theorem 1.12.
Suppose $d\in\mathbb{N}$ and $d>1$. Then for any non-negative integer $k$ with $11\leq k\leq d-$ 1, there exists a collection $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfying the following:
  • (i)
    for all $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ cannot be pierced by finitely many $k$-flats and
  • (ii)
    every strictly heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
Keller and Perles [KP22] showed that fatness assumptions can be removed for families of closed balls and point transversals.
Theorem 1.13 (Keller and Perles [KP22]: $(\aleph_{0},2)$-theorem for closed balls).
Let $\mathcal{F}$ be a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on radii). Then either $\mathcal{F}$ has a point transversal of finite size or it contains an infinite sequence of pairwise disjoint balls.
We will prove a colorful variant of the Theorem 1.13 for families of closed convex sets satisfying heterochromatic $(\aleph_{0},2)$-property. To state the result we will need some additional notations and definitions. For a set $S\subseteq\mathbb{R}^{d}$, the out-radius $R(S)$ is defined as the radius of the smallest closed ball containing $S$, and in-radius $r(S)$ of $S$ is defined as the radius of the largest closed ball contained in $C$. Condition number $\sigma\left(S\right)$ of a set $S$ is defined as $\sigma(S):=\frac{r(S)}{R(S)}$. For a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$, condition number $\sigma(\mathcal{F})$ of $\mathcal{F}$ is defined as $\sigma(\mathcal{F}):=\min\left\{\sigma(S)\;:\;S\in\mathcal{F}\right\}$.
Theorem 1.14 (Countably colorful $(\aleph_{0},2)$-theorem and bounded condition number).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying heterochromatic $(\aleph_{0},2)$-property and $\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0$. Then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many points.
As a direct corollary of the above result we get the following generalization of Theorem 1.13 by Keller and Perles.
Corollary 1.15.
Let $\mathcal{F}$ be a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F})>0$. Then either $\mathcal{F}$ has a point transversal of finite size or it contains an infinite sequence of pairwise disjoint convex sets.
Both Theorem 1.10 and Theorem 1.14 require the convex bodies to have nice shape. In Theorem 1.10, we need the convex bodies to be $(r,R)$-fat, and in Theorem 1.14 we require the condition number of the families to be bounded away from zero. It is natural to ask if one can prove a result, similar to the above mentioned theorems, for some natural families of convex bodies that are neither fat nor do they have a lower bound on the condition number. We can show the following heterochromatic $(\aleph_{0},2)$-theorem for axis-parallel boxes.
Theorem 1.16 (Heterochromatic $(\aleph_{0},2)$-theorem for boxes).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of axis-parallel boxes in $\mathbb{R}^{d}$ such that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies heterochromatic $(\aleph_{0},2)$-property. Then all but finitely many $\mathcal{F}_{n}$’s are pierced by finitely many points.
The above result is interesting because axis-parallel boxes are neither fat nor do they have a lower bound on the condition number away from zero. The following impossibility result shows that
  • Theorem 1.16 is not true if boxes are not axis-parallel.
  • Assuming the convex bodies have non-zero volumes is not sufficient to establish a $(\aleph_{0},\cdot)$-theorem.
  • A natural question related to Theorem 1.7 was, whether an $(\aleph_{0},k+2)$-theorem can be obtained for families of convex sets without the fatness assumption but the following impossibility result shows that this is not possible, even for $k=0$.
Theorem 1.17 (Impossibility result).
There exists an infinite family $\mathcal{F}$ of pairwise intersecting rectangles in $\mathbb{R}^{2}$ with bounded diameters and non-zero volumes that cannot be pierced by any finite collection of points in $\mathbb{R}^{2}$.
Remark 1.18.
Note that the proof of Theorem 1.17 can be modified for right-angled triangles in $\mathbb{R}^{2}$ with bounded diameter and non-zero volumes, see Theorem 5.1.
Organisation of the paper
  • In Section 2 we will give the definitions and notations used in this paper.
  • In Section 3 we will prove Theorem 1.10, Proposition 1.11 and Theorem 1.12.
  • Theorem 1.14 and Theorem 1.16 will be proved in Section 4.
2 Definitions and notations
We will use the following definitions and notations in this paper.
  • For any $n\in\mathbb{N}$, $[n]$ denotes the set $\left\{1,\dots,n\right\}$.
  • For any $\alpha,\beta\in\mathbb{R}$ with $\alpha\leq\beta$, $[\alpha,\beta]$ denotes the interval $[\alpha,\beta]:=\left\{x\in\mathbb{R}:\;\alpha\leq x\leq\beta\right\}$.
  • Any axis-parallel box $B$ in $\mathbb{R}^{d}$ is a closed convex set of the form $B=[\alpha_{1},\beta_{1}]\times\dots\times[\alpha_{d},\beta_{d}]$, where $\alpha_{i},\beta_{i}\in\mathbb{R}$ with $\alpha_{i}\leq\beta_{i}$, $\forall i\in[d]$.
  • Let $0<r\leq R$. A subset $A$ of $\mathbb{R}^{d}$ is called $(r,R)$-fat if $A$ contains a closed ball of radius $r$ and is contained in a closed ball of radius $R$.
  • $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$ with center at the origin $O$.
  • For any $x\in\mathbb{R}^{d}\setminus\{O\}$, $\hat{x}$ denotes the central projection of $x$ on $\mathbb{S}^{d-1}$, that is, $\hat{x}:=\frac{x}{||x||}$.
  • For any $x\in\mathbb{R}^{d}$ and a ball $B$, $\text{dist}(x,B):=\min\{\text{dist}(x,b)\;|\;b\in B\}$.
  • For any $k$ such that $10\leq k\leq d-$ 1, any $k$-dimensional affine subspace of $\mathbb{R}^{d}$ is called a $k$-flat in $\mathbb{R}^{d}$. For example any point in $\mathbb{R}^{d}$ can be considered as a 00-flat, any line in $\mathbb{R}^{d}$ as a $1$-flat, any hyperplane in $\mathbb{R}^{d}$ as $(d-1)$-flat.
  • Direction of a $k$-flat $\mathcal{K}$ is defined as follows: First $\mathcal{K}$ is translated $\mathcal{K}^{\prime}$ so that $\mathcal{K}^{\prime}$ passes through the origin. Then direction of $\mathcal{K}$ is defined as the $(k-1)$-sphere at which $\mathcal{K}^{\prime}$ intersects $\mathbb{S}^{d-1}$. This definition has been taken from Aronov, Goodman and Pollack [AGP02].
  • A $k_{1}$-flat $\mathcal{K}_{1}$ and a $k_{2}$-flat $\mathcal{K}_{2}$ are said to be parallel if one of their directions contains the other, that is, if we translate $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ to $\mathcal{K}^{\prime}_{1}$ and $\mathcal{K}^{\prime}_{2}$ respectively so that both $\mathcal{K}^{\prime}_{1}$ and $\mathcal{K}^{\prime}_{2}$ pass through the origin then either $\mathcal{K}^{\prime}_{1}\subseteq\mathcal{K}^{\prime}_{2}$ or $\mathcal{K}^{\prime}_{2}\subseteq\mathcal{K}^{\prime}_{1}$. This definition has been taken from Keller and Perles [KP22].
  • For any $x\in\mathbb{R}^{d}$ and $\varepsilon>0$, $B(x,\varepsilon)$ denotes the closed ball of radius $\varepsilon$ with centered at the origin and $B^{o}(x,\varepsilon)$ denotes the open ball of radius $\varepsilon$ with centered at the origin. $B^{o}(x,\varepsilon)$ is also considered as the (open) $\varepsilon$-neighbourhood of $x$.
  • For any $x\in\mathbb{S}^{d-1}$ and $\varepsilon>0$, an $\varepsilon$-neighbourhood of $x$ on the sphere $\mathbb{S}^{d-1}$ is $B^{o}(x,\varepsilon)\cap\mathbb{S}^{d-1}$.
  • For any $\hat{x}\in\mathbb{S}^{d-1}$ and $\varepsilon>0$ and any family of closed balls $\mathcal{F}$ in $\mathbb{R}^{d}$ the (open) $\varepsilon$-neighbourhood of $\hat{x}$ in $\mathcal{F}$ is $\mathcal{F}_{(\hat{x},\varepsilon)}=\{B(y,r)\in\mathcal{F}\;|\;\text{dist}(\hat{y},\hat{x})<\varepsilon\}$.
  • An unbounded cone $\mathcal{C}$ in $\mathbb{R}^{d}$ with vertex at the origin $O$ is a convex set such that for any $u,v\in\mathcal{C}$ and $\alpha,\beta\geq 0$, we have $\alpha u+\beta v\in\mathcal{C}$. The axis of $\mathcal{C}$ is a straight line $\ell$ passing through the vertex such that $\mathcal{C}$ has a circular symmetry about $\ell$. If we translate $\mathcal{C}$ at a point $z\in\ell$, we get the unbounded cone $\mathcal{C}^{\prime}=\{z+c\;|\;c\in\mathcal{C}\}$.
  • Limit inferior of a sequence $\left\{x_{n}\right\}_{n\in\mathbb{N}}$ of real numbers is defined as
    $\liminf_{n\to\infty}x_{n}:=\sup\left\{\inf\left\{x_{m}\;:\;m\geq n\right\}\;:\;n\in\mathbb{N}\right\}.$
  • A collection of $k$-flats $\mathcal{K}$ is said to be an $\varepsilon$-net cover of a region $P\subsetneq\mathbb{R}^{d}$ if for each point $p\in P$, there exists a $k$-flat $\kappa\in\mathcal{K}$ such that $\text{dist}(p,\kappa)<\varepsilon$.
3 Heterochromatic $(\aleph_{0},\cdot)$-theorem for fat convex sets
In this section, we will first prove Theorem 1.10, and then Proposition 1.11 and Theorem 1.12.
Let us first prove a Lemma that we will be using quite frequently. Let $\mathcal{B}$ and $\mathcal{S}$ be collections of nonempty sets in $\mathbb{R}^{d}$. For any $\mathcal{F}\subseteq\mathcal{B}$, we say $\mathcal{F}$ is finitely pierceable by $\mathcal{S}$ if there exists a finite set $S\subseteq\mathcal{S}$ such that for all $B\in\mathcal{F}$, we have $\bigcup_{A\in S}A\cap B\neq\emptyset$. A sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ $\forall n\in\mathbb{N}$ is said to have the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$ if for every heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ chosen from $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, we have a set $A\in\mathcal{S}$ such that there are $q$ distinct $B_{n}$’s that have a nonempty intersection with $A$. We say that the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, there exists an $n\in\mathbb{N}$ for which $\mathcal{F}_{n}$ is finitely pierceable by $\mathcal{S}$. We say that the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, all but finitely many $\mathcal{F}_{n}$’s are finitely pierceable by $\mathcal{S}$.
Lemma 3.1 (Heterochromatic Property Lemma).
The heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ implies the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$.
Proof.
Let the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ hold. Let $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, $\mathcal{F}_{n}\subseteq\mathcal{B}$, be a sequence that satisfies the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$. If possible, let $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ where $\{n_{i}\}_{i\in\mathbb{N}}$ is a strictly increasing sequence in $\mathbb{N}$ be such that no $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. As $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ also satisfies the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, there is an $i\in\mathbb{N}$ for which $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. But this is a contradiction, and therefore, the result follows. ∎
So, we shall be using heterochromatic $(\aleph_{0},q)$-theorem and infinite heterochromatic $(\aleph_{0},q)$-theorem interchangeably. It should be noted that for some $\mathcal{B}$ and $\mathcal{S}$ satisfying the heterochromatic $(\aleph_{0},q)$-theorem and any $N\in\mathbb{N}$, we can find a sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ that satisfy the heterochromatic $(\aleph_{0},q)$ property and for $N$ distinct values of $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is not finitely pierceable by $\mathcal{S}$. This can easily be seen by setting $\mathcal{B}$ to be the set of all unit balls in $\mathbb{R}^{d}$, $\mathcal{S}$ to be the set of all points in $\mathbb{R}^{d}$ and $\mathcal{F}_{1}=\dots=\mathcal{F}_{N}=\mathcal{B}$ and $\mathcal{F}_{n}=\{B(\mathcal{O},1)\}$ for all $n\in\mathbb{N}\setminus\{1,\dots,N\}$. See 1.10
Keller and Perles[KP22] showed that it is sufficient to prove Theorem 1.10 $(i)$ for families of closed unit balls, using the following observation.
Observation 3.2 (Keller and Perles [KP22]).
Let $d$ and $k$ be non-negative integers with $10\leq k\leq k-$ 1. Suppose $\mathcal{F}$ is a family of closed unit balls and $\mathcal{F}^{\prime}$ is a family of closed balls of radius $r$, for some $r>0$, with the same center. Then $\mathcal{F}$ is pierceable by a finite number of $k$-flats if and only if $\mathcal{F}^{\prime}$ is pierceable by a finite number of $k$-flats.
By observing that each $(r,R)$-fat closed convex set $C$ in $\mathbb{R}^{d}$, we have $B_{r}\subseteq B_{2R}$, where $B_{r}$ is a ball of radius $r$ and $B_{2R}$ is a ball of radius $2R$ with the same centre as $B_{r}$, the conclusion follows that it suffices to prove the case for $B_{2R}$’s, which we can take, without loss of generality, to be unit balls.
Before we prove Theorem 1.10 $(i)$ for closed unit balls and general $k$-flats, let’s prove some results for the special case $k=0$, that is for point transversals.
Theorem 3.3 (Heterochromatic $(\aleph_{0},2)$-theorem for fat sets).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of $(r,R)$-fat convex sets in $\mathbb{R}^{d}$ such that $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies heterochromatic $(\aleph_{0},2)$-property. Then all but finitely many $\mathcal{F}_{i}$’s are pierceable by a finite number of points.
From Observation 3.2, it is clear that to establish Theorem 3.3 proving the result for unit balls is sufficient.
Lemma 3.4.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ is not pierceable by finite number of points. Then there exists a strictly heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ that is 00-independent.
Proof.
Our claim is that given any $R>0$ and any $n\in\mathbb{N}$, there exists $B\in\mathcal{F}_{n}$ such that $\text{dist}(B,O)>R$. Since there can be at most finitely many pairwise disjoint closed unit balls in $\mathcal{F}_{n}\cap B(O,R+1)$, we can choose $B\in\mathcal{F}_{n}$ with $\text{dist}(B,O)\geq R+1$.
Now suppose for any $m\in\mathbb{N}$, $B_{1},\dots,B_{m}$ are pairwise disjoint unit balls such that $\forall i\in[m],\;B_{i}\in\mathcal{F}_{i}$. Let $\bigcup_{i\in[m]}B_{i}\subseteq B(O,R_{m})$, for some $R_{m}>0$. Then by the above claim we can choose $B_{m+1}\in\mathcal{F}_{m+1}$ such that with $\text{dist}(B_{m+1},O)\geq R_{m+1}+1$. Then $\forall i\in[m],\;B_{i}\cap B_{m+1}=\emptyset$. Thus by induction on $m$, we can get the required heterochromatic sequence. ∎
Observe that Lemma 3.4 is equivalent to the following result:
Lemma 3.5 (Contraposition of Lemma 3.4).
For each $n\in\mathbb{N}$, suppose that $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ and the collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property. Then there exists an $i\in\mathbb{N}$ such that the family $\mathcal{F}_{i}$ can be pierced by finitely many points.
Now applying Lemma 3.1, we obtain the following:
Corollary 3.6 (Heterochromatic $(\aleph_{0},2)$-theorem for closed unit balls).
For each $n\in\mathbb{N}$, suppose that $\mathcal{F}_{n}$ is a family of units balls in $\mathbb{R}^{d}$ and the collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the heterochromatic $(\aleph_{0},2)$-property. Then all but finitely many families are pierceable by finitely many points.
Now we are ready to prove Theorem 1.10 $(ii)$.
Theorem 3.7 (Restatement of Theorem 1.10 $(ii)$).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of units balls in $\mathbb{R}^{d}$ such that $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfy the weak heterochromatic $(\aleph_{0},2)$ property. Then there are at least two families $\mathcal{F}_{i},\mathcal{F}_{j},\;i\neq j$, each of which is pierceable by finite number of points.
Proof.
Using Lemma 3.4, we get that, there is at least one family $\mathcal{F}_{i}$, which is pierceable by a finite number of points. Now take any $B^{\prime}_{i}\in\mathcal{F}_{i}$ and for each $n\in\mathbb{N},\;n\neq i$, we define,
$\mathcal{F}^{\prime}_{n}=\{B\in\mathcal{F}_{n}\;|\;B\cap B^{\prime}_{i}=\emptyset\}.$
Then for each $n\in\mathbb{N},\;n\neq i,\;\mathcal{F}_{n}\setminus\mathcal{F}^{\prime}_{n}$ is pierceable by finite number of points. Then it is sufficient to show that at least one family $\mathcal{F}^{\prime}_{j},j\neq i$ is pierceable by a finite number of points. If $\mathcal{F}^{\prime}_{j}=\emptyset$, for some $j\neq i$, then we are done. Otherwise, $\forall j\in\mathbb{N},\;j\neq i,\;\mathcal{F}^{\prime}_{j}\neq\emptyset$. Now consider any heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}\setminus\{i\}}$ such that for each $n\in\mathbb{N}\setminus\{i\}$, $B_{n}\in\mathcal{F}^{\prime}_{n}$. Then $\{B_{n}\}_{n\in\mathbb{N}}$, where $B_{i}=B^{\prime}{{}_{i}}$ and $B_{n}\in\mathcal{F}^{\prime}_{n}$ for all $n\in\mathbb{N}\setminus\{i\}$, is a strict heterochromatic sequence and hence it contains at least one intersecting pair. Since for each $n\in\mathbb{N}\setminus\{i\},\;B_{n}\in\mathcal{F}^{\prime}_{n}$ implies $B_{n}\cap B^{\prime}_{i}=\emptyset$, so we must have at least one intersecting pair in $\{B_{n}\}_{n\in\mathbb{N}\setminus\{i\}}$. Using Lemma 3.4, we get that, there is at least one family $\mathcal{F}^{\prime}_{j}$, which is pierceable by a finite number of points. ∎
Figure 1: An illustration for the proof of Proposition 1.11.
Next, we shall show the tightness of the above result, that is Proposition 1.11. See 1.11
Proof.
Suppose $\mathcal{F}_{1}$ is the family containing only one unit ball $B_{1}$ in $\mathbb{R}^{2}$ with center at the origin and $\mathcal{F}_{2}$ contains the $4$ unit balls $B_{2,1},\dots,B_{2,4}$ in $\mathbb{R}^{2}$ with centers at the points $(1.5,0)$, $(-1.5,0)$, $(0,1.5)$, and $(0,-1.5)$ in $\mathbb{R}^{2}$ respectively. Then $\forall B\in\mathcal{F}_{2},\;B\cap B_{1}\neq\emptyset$ and clearly each of the families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are pierceable by a finite number of points. Now for each $n\in\mathbb{N},\;n\geq 3$, we define $\mathcal{F}_{n}$ to be the family of unit balls $\{B_{n,j}\;|\;j\in\mathbb{N}\}$, where $B_{n,j}$ is the unit ball in $\mathbb{R}^{d}$ with center at $c_{n,j}=(4j,4n),\;\forall j\in\mathbb{N}$. Then for any $m,n,i,j\in\mathbb{N},\;m,n\geq 3$ with $(m,i)\neq(n,j)$ we have $B_{m,i}\cap B_{n,j}=\emptyset$ and for each $n\in\mathbb{N},\;n\geq 3$, the family $\mathcal{F}_{n}$ is not pierceable by finite number of points. Hence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies both the properties required.
Finally, we shall prove Theorem 1.10 $(i)$ for closed unit balls. Lemma 3.1 ensures that it is sufficient to prove the following:
Theorem 3.8.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ is not pierceable by a finite number of $k$-flats, where $10\leq k\leq d-$ 1. Then there exists a $k$-independent heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of closed unit balls.
Now to prove Theorem 3.8 we will need the following lemmas.
Lemma 3.9.
Suppose $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$ such that $\mathcal{F}$ is not pierceable by a finite number of $k$-flats. Then there exists $\hat{x}\in\mathbb{S}^{d-1}$ such that $\forall\varepsilon>0,\mathcal{F}_{(\hat{x},\varepsilon)}$ is not pierceable by finite number of $k$-flats.
Proof.
The proof of this is identical to Keller and Perles’[KP22] proof for unit balls. Regardless, we add the proof here for the reader’s convenience. If possible let the statement of the lemma be false. Then $\forall\hat{y}\in\mathbb{S}^{d-1},\exists\varepsilon(\hat{y})$ for which $\mathcal{F}_{(\hat{y},\varepsilon(\hat{y}))}$ is pierceable by finite number of $k$-flats. Now $\left\{B(\hat{y},\varepsilon(\hat{y}))\;|\;\hat{y}\in\mathbb{S}^{d-1}\right\}$ is an open cover of $\mathbb{S}^{d-1}$. So, by compactness of $\mathbb{S}^{d-1},\;\exists\hat{y}_{1},\dots,\hat{y}_{r}\in\mathbb{S}^{d-1}$ such that $\mathbb{S}^{d-1}\subseteq\bigcup_{i=1}^{r}B(\hat{y}_{i},\varepsilon(\hat{y}_{i}))$ and hence $\mathcal{F}\subseteq\bigcup_{i=1}^{r}\mathcal{F}_{(\hat{y}_{i},\varepsilon(\hat{y}_{i}))}$. For each $i\in\ [r]$, since $\mathcal{F}_{(\hat{y}_{i},\varepsilon(\hat{y}_{i}))}$ is pierceable by finitely many $k$-flats, we get $\mathcal{F}$ is pierceable by finitely many $k$-flats, a contradiction. Hence the result. ∎
The following important lemma is due to Keller and Perles.
Lemma 3.10 ([KP22]).
Let $\mathcal{F}$ be a family of closed unit balls in $\mathbb{R}^{d}$, $10\leq k\leq d-$ 1 and $m\in\mathbb{N}$. If every finite subset of $\mathcal{F}$ can be pierced by $m$ $k$-flats, then $\mathcal{F}$ can be pierced by $m$ $k$-flats.
Lemma 3.11.
[…]
Proof.
Consider the space $C=B_{1}\times B_{2}\times\dots\times B_{k+1}$. Then clearly $C$ is compact. Now take any $c=(b_{1}\dots,b_{k+1})\in C$ and suppose $\kappa_{c}$ is the $k$-flat generated by $\{b_{1}\dots,b_{k+1}\}$ in $\mathbb{R}^{d}$. We define a function $f\;:\;C\rightarrow\mathbb{R}_{\geq 0}$ such that
[…]
Then $f$ is a continuous function. Since for all $\kappa\in\mathcal{K}$, direction of $\kappa$ does not contain the point $\hat{x}$, we have $f(c)>0$, for all $c\in C$. So by compactness of $C$ we get that, there exists $\varepsilon>0$ such that $f(c)\geq\varepsilon$, for all $c\in C$. ∎
Now we are ready to prove Theorem 3.8.
See 3.8
Proof.
For each $n\in\mathbb{N}$, since $\mathcal{F}_{n}$ is not pierceable by finitely many $k$-flats, by Lemma 3.9, there exists $\hat{x}_{n}\in\mathbb{S}^{d-1}$ such that $\forall\varepsilon>0,\mathcal{F}_{n(\hat{x}_{n},\varepsilon)}$ is not pierceable by finite number of $k$-flats.
Now by compactness of $\mathbb{S}^{d-1}$, $\exists\hat{x}\in\mathbb{S}^{d-1}$ such that $\{\hat{x}_{n}\}_{n\in\mathbb{N}}$ has a subsequence converging to $\hat{x}$. Without loss of generality, we may assume that the convergent subsequence itself is $\{\hat{x}_{n}\}$. Then for every $\varepsilon>0$ there exists $N(\varepsilon)\in\mathbb{N}$ such that $\forall n\geq N(\varepsilon),\;\text{dist}(\hat{x},\hat{x}_{n})<\varepsilon$.
Now we shall induct on $k$ and $d$ to prove the result. If $k=0$, then by Lemma 3.4 we get the result. So we let $k>0$. Suppose that the result holds for families of closed unit balls in $\mathbb{R}^{d-1}$ with respect to $(k-1)$-flats. We shall show that it holds for $k$-flats in $\mathbb{R}^{d}$.
Choose a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of strictly decreasing positive numbers such that $\varepsilon_{n}\to 0$ as $n\to\infty$.
Now we show that we can pick an infinite sequence $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ of unit balls such that $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ cannot be pierced by finitely many $k$-flats, and for every $i\in\mathbb{N}$, there is an $m(i)\in\mathbb{N}$ for which we have $\bigcup_{n=i}^{\infty}\mathcal{F}_{n(\hat{x}_{n},\varepsilon_{n})}\supsetneq\{B^{\prime}_{n}\}_{n=m(i)}^{\infty}$ with the property that for all $n\geq m(i)$, $B^{\prime}_{n}\cap B(\mathcal{O},i)=\emptyset$. For all $n\in\mathbb{N}$, let $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}=\{B\in\mathcal{F}_{n(\hat{x}_{n},\varepsilon_{n})}\mid B\cap B(\mathcal{O},n)=\emptyset\}$. Then clearly, $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}$ cannot be pierced by finitely many $k$-flats. Let us have $\mathcal{B}_{1},\dots,\mathcal{B}_{i-1}$ such that $\forall j\in[i-1]$, we have $\mathcal{B}_{j}\subsetneq\mathbb{F}^{\prime}_{j(\hat{x}_{j},\varepsilon_{j})}$ and $\mathcal{B}_{j}$ is finite. Then there is an $M(i-1)\in\mathbb{N}$ such that at least $M(i-1)$ $k$-flats are needed to pierce every set in $\bigcup_{j=1}^{i-1}\mathcal{B}_{j}$. We show that there is a finite set $\mathcal{B}_{i}\subsetneq\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ such that at least $M(i)\geq M(i-1)+1$ $k$-flats are needed to pierce every set in $\bigcup_{j=1}^{i}\mathcal{B}_{j}$. If not, then every finite subset of the set $\mathcal{B}_{1}\cup\dots\cup\mathcal{B}_{i-1}\cup\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$, and therefore $\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ has a $k$-transversal of size $M(i-1)$. Then by Lemma 3.10, $\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ is pierceable by $m(i-1)$ $k$-flats, which is a contradiction. Therefore, we get a finite set $\mathcal{B}_{i}\subsetneq\mathcal{F}^{\prime}_{i(\hat{x}_{i},\varepsilon_{i})}$ for which $\bigcup_{j=1}^{i}\mathcal{B}_{j}$ cannot be pierced by fewer than $M(i)$ $k$-flats. Continuing this way, we get our sequence $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}=\bigcup_{j=1}^{\infty}\mathcal{B}_{j}$ which is not pierceable by finitely many $k$-flats. Let =$\ell=\{R\hat{x}\mid R\geq 0\}$. Let $S$ denote the $(d-1)$-flat through $\mathcal{O}$ and perpendicular to $\ell$, and let for every $B^{\prime}_{n}$, $\pi(B^{\prime}_{n})$ denote the projection of $B^{\prime}_{n}$ onto $S$. The sequence $\{\pi(B^{\prime}_{n})\}_{n\in\mathbb{N}}$ then does not have a finite $(k-1)$-transversal, because then we would have finitely many $k$-flats parallel to $\ell$ piercing $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$. So, there is a subsequence of $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ whose projection onto $S$ is $(k-1)$-independent. Without loss of generality, let that subsequence be $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$. For any $k+1$ sets $B_{1},\dots,B_{k+1}$ picked from $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$, there is no $k$-flat parallel to $\ell$ passing through every $B_{i}$, $11\leq i\leq k+$ 1. That is because no $(k-1)$-flat passes through the projections of $B_{1},\dots,B_{k+1}$ on $S$, since the intersection of a $k$-flat perpendicular to $\ell$ with $S$ is a $(k-1)$-flat on $S$. Now let $B_{1},\dots,B_{m}$ be picked from $\{B^{\prime}_{n}\}_{n\in\mathbb{N}}$ with $B_{j}\in\mathcal{B}_{n_{j}}$, $n_{j}\neq n_{j}^{\prime}$ for $j\neq j^{\prime}$, $m\geq k+1$, such that no $k$-flat intersects more than $k+1$ distinct $B_{i}$’s and no $k$-flat perpendicular to $\ell$ passes through $k+1$ distinct $B_{i}$’s. Then $B_{1},\dots,B_{m}$ satisfy the conditions of Lemma 3.11, and therefore, there exists an $\varepsilon>0$ such that the direction of every $k$-flat intersecting all of $B_{1},\dots,B_{m}$ is at least $\varepsilon$-distance apart from $\hat{x}$. There is an $i_{1}\in\mathbb{N}$ such that for all $n\geq i_{1}$, $B(\hat{x}_{n},\varepsilon_{n})\subset B(\hat{x},\frac{\varepsilon}{2})$. Corresponding to this $i_{1}$, there is an $m(i_{1})\in\mathbb{N}$ such that $\bigcup_{n=i_{1}}^{\infty}\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}\supsetneq\{B^{\prime}_{n}\}_{n=m(i_{1})}^{\infty}$. Also, let $i_{2}=\max\{n_{1},\dots,n_{m}\}$, where $B_{j}\in\mathcal{B}_{n_{j}}$. Now, it is easy to check that there is a positive integer $i_{3}>0$ for which no $k$-flat that passes through any $k+1$ distinct $B_{i}$’s chosen from $B_{1},\dots,B_{m}$ intersects any sets in $\mathcal{F}^{\prime}_{n(\hat{x}_{n},\varepsilon_{n})}$ whenever $n>i_{3}$, since the angle that any such $k$-flat makes with $\ell$ is bounded by positive numbers. Set $N=1+\max\{i_{1},i_{2},i_{3}\}$, and set $B_{m+1}$ to be any set from $\{B^{\prime}_{n}\}_{n=m(N)}^{\infty}$. Continuing this way, we obtain our heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$.
It is quite intuitive from the proof technique of Theorem 3.8 that for higher order transversals, that is for $k$-transversals with $k>0$, the existence of the strictly heterochromatic sequence of balls is not necessarily ensured if the radius of the balls tends to zero. Actually we can show that even if all the balls have a fixed radius, then also the existence of the strictly heterochromatic sequence of balls is not guaranteed.
See 1.12
Figure 2: Dotted region denotes $A_{\{j_{1},j_{2}\}}$
Proof.
Suppose $\mathcal{I}^{\prime}$ is the collection of all multisets of size $(k+1)$ from $\mathbb{N}$, i.e, each $i\in\mathcal{I}^{\prime}$ is of the form $i=\{i_{1},\dots,i_{k+1}\}$ with each $i_{j}\in\mathbb{N}$ and all $i_{j}$’s are not necessarily distinct. Next suppose that $\mathcal{I}=[k+1]\cup\mathcal{I}^{\prime}$, i.e, $i\in\mathcal{I}$ implies either $i$ is a natural number lying in $[k+1]$ or $i$ is a multiset of size $(k+1)$ from $\mathbb{N}$. Then clearly $\mathcal{I}$ has a bijection with $\mathbb{N}$.
Now we take an infinite sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of closed unit balls in $\mathbb{R}^{d}$ such that no $(k+2)$ balls from the sequence are pierceable by a single $k$-flat.
For each $\ell\in[k+1]$, define $\mathcal{F}_{\ell}$ to be the family $\{B_{n}\}_{n\in\mathbb{N}}$. And for any $i\in\mathcal{I}$ with the form $i=\{{j_{1}},\dots,j_{k+1}\}$, we define the region
$A_{i}:=\{x\in\mathbb{R}^{d}\;|\;\text{there is a $k$-flat passing through $x$ that pierces the balls }B_{j_{1}},\dots,B_{j_{k+1}}\},$
see Figure 2. Next we define the family $F_{i}$ to be the collection that we obtain by tightly packing the region $A_{i}$ with closed unit balls. Then $\mathcal{F}_{i}$ is not pierceable by finitely many $k$-flats. Thus for each $i\in\mathcal{I}$, we define $\mathcal{F}_{i}$ to be a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{i}$ is not pierceable by a finite number of $k$-flats.
Now we choose any $k+1$ balls from the families $\mathcal{F}_{1},\dots,\mathcal{F}_{k+1}$, say, $B_{j_{1}},\dots,B_{j_{k+1}}$, where $B_{j_{\ell}}\in\mathcal{F}_{\ell},\forall\ell\in[k+1]$ and consider the $(k+1)$-tuple $i=\{{j_{1}},\dots,j_{k+1}\}$. Then $i\in\mathcal{I}$. Then by construction of the family $\mathcal{F}_{i}$, we get that each $B\in\mathcal{F}_{i}$ is pierceable by a $k$-flat that pierces all the $(k+1)$ balls $B_{j_{1}},\dots,B_{j_{k+1}}$. This proves the first part of the theorem. ∎
4 Heterochromatic $(\aleph_{0},2)$-Theorem for convex sets
In this section, we will prove Theorem 1.14 and Theorem 1.16. To prove Theorem 1.14, we will begin by showing that it is sufficient to establish the following result.
Theorem 4.1.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F}_{n})=\sigma_{n}$ such that $\sigma_{n}\geq\sigma_{0}>0$ and $\mathcal{F}_{n}$ is not pierceable by finite number of points. Then $\exists$ a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ that is 00-independent.
Observation 4.2.
Theorem 4.1 implies Theorem 1.14.
Proof.
Suppose Theorem 4.1 holds and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying heterochromatic $(\aleph_{0},2)$-property and
$\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0.$
Then there exists $\sigma_{0}>0$ and a subsequence $\left\{\sigma_{n_{i}}\right\}_{i\in\mathbb{N}}$ such that $\forall i\in\mathbb{N},\;\sigma_{n_{i}}>\sigma_{0}$ and applying Theorem 4.1 on the families $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$, we get that there exists $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ is pierceable by finite number of points. Without loss of generality, we may assume that $\mathcal{F}_{n_{1}}$ is pierceable by a finite number of points.
Again applying the same argument over the families $\{\mathcal{F}_{n_{i}}\}_{i>1}$, we get that there exists $i\in\mathbb{N}$ with $i>1$ such that $\mathcal{F}_{n_{i}}$ is pierceable by finite number of points. Continuing this we get that there are infinitely many $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ we get that there exists $i\in\mathbb{N}$ such that $\mathcal{F}_{n_{i}}$ is pierceable by a finite number of points. ∎
Now to prove Theorem 4.1 we will need the following technical lemma:
Lemma 4.3.
[…]
Proof.
For each $x\in\mathcal{B}$, suppose $\mathcal{F}_{x}=\{S\in\mathcal{F}\;|\;\text{dist}(x,S)<\delta_{x}\}$. Since for each $S\in\mathcal{F}_{x}$ either $r(S)\geq\delta_{x}\sigma_{0}$ or $x\in S$, $\mathcal{F}_{x}$ is pierceable by finite number of points, say by $t_{x}=t_{x}(\delta_{x})$ points. Now consider the open cover $\{B^{o}(x,\delta_{x})\;|\;x\in\mathcal{B}\}$ of $\mathcal{B}$. Then by the compactness of $\mathcal{B}$, this open cover has a finite subcover, that is, $\exists n\in\mathbb{N}$ such that $\mathcal{B}\subseteq\bigcup_{i=1}^{n}B^{o}(x_{i},\delta_{x_{i}})$. Now we claim that $\mathcal{F}\subseteq\bigcup_{i=1}^{n}\mathcal{F}_{x_{i}}$. Take any $S\in\mathcal{F}$. Since $S\cap\mathcal{B}\neq\emptyset$, $\exists x\in\mathcal{B}\cap S$. Now $\exists i\in[n]$ such that $x\in B^{o}(x_{i},\delta_{x_{i}})$ and hence $\text{dist}(S,x_{i})<\delta_{x_{i}}\implies S\in\mathcal{F}_{x_{i}}$. So $\mathcal{F}\subseteq\bigcup_{i=1}^{n}\mathcal{F}_{x_{i}}$. For each $i\in[n]$, since $\mathcal{F}_{x_{i}}$ is pierceable by $t_{x_{i}}$ points, so we get that $\mathcal{F}$ is pierceable by $\sum_{i=1}^{n}t_{x_{i}}$ points. ∎
Since for each $S\in\mathcal{F}$, $\frac{r(S)}{R(S)}\geq\sigma_{0}$, the following corollary directly from the above lemma.
Corollary 4.4.
[…]
Now we are ready to prove Theorem 4.1. See 4.1
Proof.
For any $m\in\mathbb{N}$, suppose $S_{1},\dots,S_{m}$ are pairwise disjoint sets such that each $S_{i}$ comes from a different family. Without loss of generality, we assume that $\forall i\in[m],\;S_{i}\in\mathcal{F}_{i}$. We shall show that either $\exists S_{m+1}\in\mathcal{F}_{p}$, for some $p>m$, such that $\forall i\in[m],\;S_{i}\cap S_{m+1}=\emptyset$ or we can identify an infinite sequence satisfying all the required conditions.
In the first case, that is, if there exists $S_{m+1}\in\mathcal{F}_{p}$, for some $p>m$, such that $\forall i\in[m],\;S_{i}\cap S_{m+1}=\emptyset$, then there is nothing to prove. Otherwise, $\forall p>m$ and $\forall S\in\mathcal{F}_{p}$, there exists $i\in[m]$ such that $S\cap S_{i}\neq\emptyset$. Then for each $p>m,\;\exists i_{p}\in[m]$ such that infinitely many members of $\mathcal{F}_{p}$ meet $S_{i_{p}}$. So $\exists\ell\in[m]$ and an infinite subset $\mathcal{J}\subseteq\{p\in\mathbb{N}\;|\;p>m\}$ such that for each $j\in\mathcal{J}$, there exists an infinite sequence $\left\{S^{(j)}_{i_{n}}\right\}_{n\in\mathbb{N}}$ satisfying the following:
  • $\forall n\in\mathbb{N},\;S^{(j)}_{i_{n}}\in\mathcal{F}_{j}$ and
  • $\forall j\in\mathcal{J},\;\forall n\in\mathbb{N},\;S^{(j)}_{i_{n}}\cap S_{\ell}\neq\emptyset$.
Without loss of generality we assume that $\forall j\in\mathcal{J},\;\mathcal{F}_{j}=\left\{S^{(j)}_{i_{n}}\right\}_{n\in\mathbb{N}}$ and $\ell=1$.
For each $j\in\mathcal{J}$, since $\mathcal{F}_{j}$ is not pierceable by finite number of points and $\forall S\in\mathcal{F}_{j},\;S\cap S_{1}\neq\emptyset$, so using Corollary 4.4, we get the following: $\forall j\in\mathcal{J},\;\exists x_{j}\in S_{1}$ such that $\forall\delta>0\;\exists S_{j}\in\mathcal{F}_{j}$ satisfying: (a) $\text{dist}(S_{j},x_{j})<\delta$, (b) out-radius of the set $S_{j}$, $R(S_{j})<\delta$ and (c) $x_{j}\not\in S_{j}$.
Then by compactness of $S_{1}$, $\exists x\in S_{1}$ such that $\{x_{j}\}_{j\in\mathcal{J}}$ has a subsequence converging to $x$. Without loss of generality, we assume that the whole sequence $\{x_{j}\}_{j\in\mathcal{J}}$ converges to $x$. We also assume that $\mathcal{J}=\{j_{n}\;|\;n\in\mathbb{N}\}$ and $j_{p}<j_{q}\text{ if and only if }p<q$ and $\forall j\in\mathcal{J},\;x_{j}\neq x$, since the case where $x_{j}=x$ for infinitely many values of $j$ is trivially inferred from Corollary 4.4. We shall now show that there exists an infinite sequence of sets satisfying all the required properties.
Suppose $\Delta_{1}:=\text{dist}(x_{j_{1}},x)\;\mbox{and}\;\delta_{1}:=\frac{\Delta_{1}}{4}.$ Then $\exists S_{j_{1}}\in\mathcal{F}_{j_{1}}$ such that $\text{dist}(S_{j_{1}},x_{j_{1}})<\delta_{1}$, and out-radius of $S_{j_{1}}$, $R(S_{j_{1}})<\delta_{1}$ and $x_{j_{1}}\not\in S_{j_{1}}$. Then we get that, $x\not\in S_{j_{1}}$, that is, $\text{dist}(x,S_{j_{1}})>0$. For otherwise, $x\in S_{j_{1}},\text{ that implies }\text{dist}(c,x)<R(S_{j_{1}})<\delta_{1},$ where $S\subseteq B(c,R(S_{j_{1}}))$. Then
$\Delta_{1}$ $=\text{dist}(x,x_{j_{1}})\leq\text{dist}(x,c)+\text{dist}(c,x_{j_{1}})\leq\text{dist}(x,c)+\text{dist}(S_{j_{1}},x_{j_{1}})+R(S_{j_{1}})<3\delta_{1},$
that is, $\Delta_{1}<3\delta_{1}$, a contradiction.
Since the sequence $\{x_{j}\}_{j\in\mathcal{J}}$ converges to $x$, $\exists j_{p}\in\mathcal{J}$ such that $\text{dist}(x,x_{j_{p}})<\text{dist}(x,S_{j_{1}})/2.$ Without loss of generality, we assume that $p=2$. Suppose $\Delta_{2}:=\text{dist}(x,x_{j_{2}})\;\mbox{and}\;\delta_{2}:=\frac{\Delta_{2}}{4}.$ Then $\exists S_{j_{2}}\in\mathcal{F}_{j_{2}}$ such that $\text{dist}(S_{j_{2}},x_{j_{2}})<\delta_{2}$, out-radius of $S_{j_{2}}$, $R(S_{j_{2}})<\delta_{2}$ and $x_{j_{2}}\not\in S_{j_{2}}$. Then similarly we can show that $x\not\in S_{j_{2}}$, that is, $\text{dist}(x,S_{j_{2}})>0$. Also, we have $S_{j_{1}}\cap S_{j_{2}}=\emptyset$. For otherwise, $\exists z\in S_{j_{1}}\cap S_{j_{2}}$. Then we have,
$2\Delta_{2}=2\text{dist}(x,x_{j_{2}})$ $<\text{dist}(x,S_{j_{1}})$
$=\min\limits_{s\in S_{j_{1}}}\text{dist}(x,s)$
$\leq\text{dist}(x,z)$
$\leq\text{dist}(x,x_{j_{2}})+\text{dist}(x_{j_{2}},z)$
$\leq\text{dist}(x,x_{j_{2}})+\text{dist}(x_{j_{2}},S_{j_{2}})+R(S_{j_{2}})$
$<\Delta_{2}+2\delta_{2}=3\Delta_{2}/2,$
that is, $\Delta_{2}<3\Delta_{2}/4$, a contradiction. Continuing this procedure, we will get an infinite heterochromatic sequence of pairwise disjoint sets. ∎
Observe that if $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$, then the condition number of $\mathcal{F}$ is $1$. This gives us the following corollary.
Corollary 4.5 (Heterochromatic $(\aleph_{0},2)$-Theorem for closed balls).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on the radii) and $\mathcal{F}_{n}$ is not pierceable by a finite number of points. Then there exists a heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of pairwise disjoint balls.
Again the following proposition is a special case of the above theorem.
Proposition 4.6 (Heterochromatic $(\aleph_{0},2)$-theorem for intervals).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed intervals in $\mathbb{R}$ such that $\mathcal{F}_{n}$ is not pierceable by a finite number of points. Then there exists a heterochromatic sequence $\{I_{n}\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of pairwise disjoint intervals.
We now prove Theorem 1.16 using the above proposition.
See 1.16
Proof.
Suppose for any box $B$ and $i\in[d]$, $\pi_{i}(B)$ is the projection of $B$ on the $i$-th axis and $\mathcal{F}_{n,i}=\left\{\pi_{i}(B)\;|\;B\in\mathcal{F}_{n}\right\}$, that is, $\mathcal{F}_{n,i}$ is the collection of closed intervals on real line obtained by projecting the members of $\mathcal{F}_{n}$ on $i$-th axis. Since for each $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is not pierceable by a finite number of points, there exists $i\in[d]$ such that $\mathcal{F}_{n,i}$ is not pierceable by a finite number of points. So there exists an $i\in[d]$ and an infinite subset $\mathcal{I}\subset\mathbb{N}$ such that $\forall n\in\mathcal{I}$, $\mathcal{F}_{n,i}$ is not pierceable by a finite number of points. Then, applying Proposition 4.6, we get that $\exists$ a heterochromatic infinite sequence of pairwise disjoint intervals $\left\{\pi_{i}(B_{n})\right\}_{n\in\mathcal{I}^{\prime}}$ such that $\mathcal{I}^{\prime}\subseteq\mathcal{I}$ is infinite. Then the corresponding boxes $B_{n}\in\mathcal{F}_{n},n\in\mathcal{I}^{\prime}$ form a desired sequence. ∎
5 Impossibilities of $(\aleph_{0},2)$-Theorem for rectangles and triangles
In this section, we will prove theorem 1.17.
See 1.17
We will first fix some notations that will be used in the proof of Theorem 1.17. Let’s fix some notations:
  • We will denote by $\mathbb{S}^{1}$ as the unit circle center at the origin in $\mathbb{R}^{2}$.
  • For all $n\in\mathbb{N}$, we will denote by $\theta_{n}$ the angle ${\pi}/{2^{n}}$.
  • Let $a=(0,1)$, $b=(1,0)$, $c=(1,1)$, and $u_{n}:=(\cos\theta_{n},\sin\theta_{n})$ for all $n\in\mathbb{N}$.
  • For all $n\in\mathbb{N}$, let $a_{n}$ and $b_{n}$ be the intersection points of the tangent to the circle $\mathbb{S}^{1}$ at the point $u_{n}$ with the lines $x_{2}=1$ and $x_{1}=1$ respectively.
  • Let :=$\mathcal{R}:=\left\{(\overline{x}_{1},\overline{x}_{2})\in\mathbb{R}^{2}\;\mid\;0\leq\overline{x}_{1}\leq 1\;\mbox{and}\;0\leq\overline{x}_{2}\leq 1\right\}$.
  • For all $n\in\mathbb{N}$, $s_{n}$ denotes the closed line segment joining the points $a_{n}$ and $b_{n}$. Let
    $B^{(i)}_{n}:=\left\{\overline{x}+\lambda u_{n}\;\mid\;\overline{x}\in s_{n},\,\lambda\in\left[0,1/2^{i}\right]\right\},$
    and $\mathcal{F}_{i}:=\left\{B_{n}^{(i)}\;\big{|}\;n\in\mathbb{N}\right\}$. See Figure 3.
Figure 3: The first figure shows the constructed line segments $s_{n}$, and the second figure shows the rectangles $B_{n}^{(i)}$.
Consider the family of rectangles
$\mathcal{F}:=\bigcup_{i\in\mathbb{N}}\mathcal{F}_{i}.$
Observe that for all $m$ and $n$ in $\mathbb{N}$, $s_{m}\cap s_{n}\neq\emptyset$, see Figure 4. Therefore, $B^{(i)}_{n}\cap B^{(j)}_{m}\neq\emptyset$. We will now show that $\mathcal{F}$ cannot be pierced by a finite set of points. To reach a contradiction suppose there exists a finite set $C\subsetneq\mathbb{R}^{2}$ such that for all $B_{n}^{(i)}\cap C\neq\emptyset$. Without loss of generality we can assume that the points $a=(0,1)$, $b=(1,0)$, and $c=(1,1)$ are in the set $C$.
First, we want to show that it will be sufficient to consider the case where
$C\subseteq\mathcal{R}\setminus B^{o}(O,1).$
Observe that by the construction of $\mathcal{F}$ if $C$ contains a point $p=(p_{1},p_{2})$ with $p_{1}<0$ or $p_{2}<0$ then $p$ does not hit any rectangles in $\mathcal{F}$. Therefore, we can assume that for all $p=(p_{1},p_{2})\in C$ we have $p_{1}\geq 0$ and $p_{2}\geq 0$. Now, if $C$ contains a point $p=(p_{1},p_{2})$ with $p_{1}>1$ then $C\cup\{\widetilde{p}\}\setminus\{p\}$ where $\widetilde{p}=(1,p_{2})$ is a point transversal of $\mathcal{F}$. Similarly, if $C$ contains a point $q=(q_{1},q_{2})$ with $q_{2}>1$ then $C\cup\{\widetilde{q}\}\setminus\{q\}$ where $\widetilde{q}=(q_{1},1)$ is also a point transversal of $\mathcal{F}$. Observe that none of the points in $B^{o}(O,1)$ can hit any rectangles in $\mathcal{F}$. Therefore, combining all the above observations we can safely assume that $C\subseteq\mathcal{R}\setminus B^{o}(O,1)$.
Figure 4: The first figure shows that any two segments always intersect, and the second figure shows how the rectangle $B_{n_{0}}^{(i_{0})}$ is disjoint from the set $C$.
Let $C_{1}$ be the subset of $C$ obtained by intersecting $C$ with the line $\ell_{1}:x_{1}-1=0$, and $C_{2}:=C\setminus C_{1}$. Also, let
$\delta:=\min_{q\in C_{2}}\text{dist}(q,\ell_{1}),\;\mbox{and}\;\lambda:=\min_{p\in C_{1}\setminus\{b\}}\|p-b\|.$
Observe that as
$\lim_{n\to\infty}a_{n}=(1,1)\;\mbox{and}\;\lim_{n\to\infty}b_{n}=(1,0),$
therefore there exists $n_{0}\in\mathbb{N}$ such that
$\|a_{n_{0}}-c\|<\delta\;\mbox{and}\;\|b_{n_{0}}-b\|<\lambda.$
See Figure 4.
Figure 5: The figure shows the triangles $T^{(i)}_{n}$ that can be used to establish Theorem 5.1.
Let $q_{1}$ be the point in the set $C_{1}\setminus\{b\}$ closest to $b$, and let $\ell_{n_{0}}$ be the line tangent to the circle $\mathbb{S}^{1}$ at the point $u_{n}$. Note that there exits $i_{0}\in\mathbb{N}$ such that
$1/2^{i_{0}}<\text{dist}(q_{1},\ell_{n_{0}}),$
see Figure 4. Now observe that the rectange $B_{n_{0}}^{(i_{0})}$ and the set $C$ is disjoint. Therefore, $C$ is not piercing $\mathcal{F}$ and we have reached a contradiction. This completes the proof of Theorem 1.17.
Observe that the construction can also be done along similar lines for non-axis parallel right-angled triangles in $\mathbb{R}^{2}$, and therefore we can prove the following result.
Theorem 5.1.
There exists a family $\mathcal{F}$ of pairwise intersecting right-angled triangles in $\mathbb{R}^{2}$ that cannot be pierced by any finite collection of points in $\mathbb{R}^{2}$. Note that the triangles in $\mathcal{F}$ have bounded diameters and non-zero volumes.