### v5

# Heterochromatic Geometric Transversals of Convex sets

Abstract
An infinite sequence of sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of families $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$, if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. In this paper, we have proved that if for each $n\in\mathbb{N},\mathcal{F}_{n}$ is a family of nicely shaped convex sets in $\mathbb{R}^{d}$ such that each heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ contains at least one $k+2$-tuple that can be pierced by a single $k$-flat ($k$-dimensional affine space) then all but finitely many $\mathcal{F}_{n}$’s can be pierced by finitely many $k$-flats. This result can be considered as a countably colorful generalizations of the $(\aleph_{0},k+2)$-theorem proved by Keller and Perles (Symposium on Computational Geometry 2022). We have also established the tightness of our result by proving a number of no-go theorems.
1 Introduction
The following theorem due to Helly [Hel23], proved more than 100 years ago, has been one of the fundamental results in discrete and convex geometry with numerous extensions and applications.
Theorem 1.1 (Helly [Hel23]).
Let $\mathcal{F}$ be an infinite family of compact convex sets in $\mathbb{R}^{d}$ such that the intersection of every $d+1$ sets from $\mathcal{F}$ is non-empty then the intersection of all the sets in $\mathcal{F}$ is non-empty.
For a comprehensive introduction to Helly’s Theorem and related topics see the recent surveys by Amenta, Loera, and Soberón [ADLS17], and Bárány and Kalai [BK22]. Unless otherwise stated explicitly, whenever we talk about a family of convex sets we won’t assume anything about the cardinality of the family, that is, a family of convex sets can be a finite collection or an infinite collection.
A $k$-flat $H$ is said to hit or pierce a set $F\subseteq\mathbb{R}^{d}$ if $F\cap H\neq\emptyset$. A set $\mathcal{T}$ of $k$-flats is called a $k$-transversal of a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ if for every $F\in\mathcal{F}$ there exists a $k$-flat in $\mathcal{T}$ that pierces $F$. Additionally, if $k=0$ or $k=1$ or $k=d-1$ then $\mathcal{T}$ is called point transversal, line transversal and hyperplane transversal of $\mathcal{F}$ respectively. Also, we say a family $\mathcal{F}$ has a $k$-transversal of size $m$ (or, at most $m$) if there exists a $k$-transversal $\mathcal{T}$ of $\mathcal{F}$ and $|\mathcal{T}|=m$ (or, $|\mathcal{T}|\leq m$). Generalizations of Helly’s Theorem to general $k$-transversals, for $11\leq k\leq d-$ 1, has been an active area of research, see the survey by Holmsen and Wenger [HW17]. Santaló [San40] showed that there can be no analog of Helly Theorem for arbitrary families of convex sets and line transversals. Because of Santaló’s impossibility result, the follow-up works on $k$-transversals are for families of convex sets with additional properties, for example, families of unit balls, etc. Hadwiger [Had56] proved the following Helly-type result for line transversal.
Theorem 1.2 (Hadwiger [Had56]).
Let ${C}$ be a compact convex set in $\mathbb{R}^{d}$ with a non-empty interior, and $\mathcal{F}$ be a countably infinite collection of disjoint convex bodies congruent to $\mathcal{C}$. If every subset of $\mathcal{F}$ of size $d+1$ has a line transversal of size one then the whole family $\mathcal{F}$ has a line transversal of size one.
Danzer, Grünbaum and Klee [DGK63] showed that the ”congruent” assumption in Theorem 1.2 can be weakened as long as the sets are of bounded diameter, that is, there exists $R>0$ such that the diameter of all the convex sets in the family is at most $R$.
Theorem 1.3 (Danzer, Grünbaum, and Klee [DGK63]).
Suppose $R>0$ and $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ with diameters at most $R$ and $\bigcup_{C\in\mathcal{F}}C$ is unbounded. If every subset of $\mathcal{F}$ of size $d+1$ has a line transversal of size one then the whole family $\mathcal{F}$ has a line transversal of size one.
A family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is said to satisfy $(p,q)$-property with respect to $k$-flats, for $10\leq k\leq d-$ 1, if every subset of $\mathcal{F}$ of size $p$ contains $q$ sets that can be pierced by a single $k$-flat. Hadwiger and Debrunner [HD57] were the first to introduce $(p,q)$-property, and they proved that for all $p\geq q\geq d+1$ satisfying $p(d-1)<(q-1)d$, if $\mathcal{F}$ is a family of compact convex sets in $\mathbb{R}^{d}$ satisfying $(p,q)$-property with respect to points then $\mathcal{F}$ has a point transversal of size at most $p-q+1$. Hadwiger and Debrunner [HD57] asked the following question that remained open for $35$-years.
Question 1.4 (Hadwiger and Debrunner $(p,q)$-problem [HD57]).
For all $p\geq q\geq d+1$, does there exists a constant $c(p,q,d)>0$ such that if a family $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{d}$ satisfy $(p,q)$-property with respect points then $\mathcal{F}$ has a point transversal of size at most $c(p,q,d)$?
Alon and Kleitman [AK92] proved the following fundamental theorem and since then this area has been one of the most active areas of research in Discrete Geometry.
Theorem 1.5 (Alon-Kleitman [AK92]: $(p,q)$-theorem).
For any three natural numbers $p\geq q\geq d+1$, $\exists c=c(p,q,d)$ such that if $\mathcal{F}$ is a collection of compact convex sets satisfying the $(p,q)$-property in $\mathbb{R}^{d}$ then there exists a point transversal for $\mathcal{F}$ of size at most $c$.
Alon and Kalai [AK95] later proved the $(p,q)$-theorem for hyperplane transversal.
Theorem 1.6 (Alon-Kalai [AK95]: $(p,q)$-theorem for hyperplane transversal).
For any three natural numbers $p\geq q\geq d+1$, $\exists c^{\prime}=c^{\prime}(p,q,d)$ such that if $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ satisfying the $(p,q)$-property with respect to hyperplanes then there exists a hyperplane transversal for $\mathcal{F}$ of size at most $c^{\prime}$.
Unlike the case of hyperplane transversal Alon, Kalai, Matoušek, and Meshulam [AKMM02] gave an explicit construction showing the impossibility of getting a $(p,q)$-theorem, like Theorem 1.6, for $k$-transversal when $10<k<d-$ 1.
Keller and Perles introduced an interesting variation of the $(p,q)$-property. A family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is said to satisfy $(\aleph_{0},q(k))$-property with respect to $k$-flats if among every infinite sequence of sets from $\mathcal{F}$ there exist $q(k)$ sets such that these $q(k)$ sets can be pierced by a single $k$-flat. Keller and Perles [KP22] proved the following important theorem for families of $(r,R)$-fat convex sets satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats. Note that a convex set $C$ is $(r,R)$-fat if there exist two closed balls $B(c_{1},r)$ and $B(c_{2},R)$ with $B(c_{1},r)\subseteq C\subseteq B(c_{2},R).$
Theorem 1.7 (Keller and Perles [KP22]: $(\aleph_{0},k+2)$-theorem for $k$-transversal).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ is an infinite family of $(r,R)$-fat convex sets, with $0<r<R<\infty$, satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats, then the whole family $\mathcal{F}$ has a finite size $k$-transversal.
Observe that for $k$-transversal $q(k)=k+2$ is optimal since any $k+1$ nonempty subset of $\mathbb{R}^{d}$ can be pierced by a single $k$-flat. Keller and Perles proved a similar result for point transversal of families of closed balls (with no restriction on radii).
Theorem 1.8 (Keller and Perles [KP22]: $(\aleph_{0},2)$-Theorem for point transversal).
Suppose $\mathcal{F}$ is a family of closed balls in $\mathbb{R}^{d}$ (with no restriction on the radii) satisfying the $\left(\aleph_{0},2\right)$-property for points. Then $\mathcal{F}$ has a finite-size point transversal.
To complement the above result, Keller and Perles also gave an explicit example showing that the compactness assumption in Theorem 1.7 cannot be weakened. Additionally, Keller and Perles observed that, unlike $(p,q)$-property, for all $10\leq k\leq d-$ 1 there does not exist a constant $c(d,k)>0$ such that if $\mathcal{F}$ is a family of unit balls in $\mathbb{R}^{d}$ satisfying $(\aleph_{0},k+2)$-property with respect to $k$-flats then $\mathcal{F}$ has a $k$-transversal of size at most $c(d,k)$. In an updated version [KP23] Keller and Perles have extended Theorem 1.7 to families of near-balls. A collection $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$ is called a family of near-balls if there exists a constant $\alpha>0$ such that for all $C\in\mathcal{F}$, there exist closed balls $B(p_{C},r_{C})$ and $B(p_{C},R_{C})$ satisfying the following two conditions:
$B(p_{C},r_{C})\subseteq C\subseteq B(p_{C},R_{C}),~{}\mbox{and}$ (1)
$R_{C}\leq\min\left\{r_{C}+\alpha,\alpha r_{C}\right\}.$ (2)
Theorem 1.9 (Keller and Perles [KP23]: $(\aleph_{0},k+2)$-Theorem for near balls).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ be a collection of compact near-balls satisfying the $(\aleph_{0},k+2)$-property for $k$-flats. Then $\mathcal{F}$ has a finite-size $k$-transversal.
Observe that the above result directly gives the following corollary.
Corollary 1.10 (Keller and Perles [KP23]: $(\aleph_{0},k+2)$-Theorem for closed balls).
Let $d$ be a natural number and $k$ be a non-negative integer with $10\leq k\leq d-$ 1. Suppose $\mathcal{F}$ be a collection of closed balls in $\mathbb{R}^{d}$ satisfying the $(\aleph_{0},k+2)$-property for $k$-flats. Then $\mathcal{F}$ has a finite-size $k$-transversal.
Note that Corollary 1.10 is a generalization of Theorem 1.7. Jung and Pálvölgyi [JP23] recently developed a new framework using which one can lift $(p,q)$-type theorems for finite families to prove new $(\aleph_{0},q)$-Theorems for infinite families, and also provided new proofs of already existing results like Corollary 1.10.
Given $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ families of convex sets in $\mathbb{R}^{d}$, we call a collection $B_{1},\dots,B_{p}$ of $p$ convex sets colorful with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ if for all $i\in\left\{1,\dots,p\right\}$ we have $B_{i}\in\mathcal{F}_{i}$. We say $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ satisfies colorful $(p,q)$-property with respect to $k$-flats if every colorful collection of $p$ sets with respect to $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ contains $q$ sets that can be pierced by a single $k$-flat. Lovász observed the following interesting variant of Helly’s theorem, called the Colorful Helly’s theorem, in 1974, though the proof first appeared in a paper on Carathéodory’s theorem by Bárány [Bár82].
Theorem 1.11 (Bárány [Bár82]: Colorful Helly’s theorem).
Suppose $\mathcal{F}_{1},\dots,\mathcal{F}_{d+1}$ are families of convex sets in $\mathbb{R}^{d}$ satisfying the colorful $(d+1,d+1)$-property with respect to points. Then $\exists\,i\in\left\{1,\dots,d+1\right\}$ and a point in $\mathbb{R}^{d}$ that pierces all the sets in $\mathcal{F}_{i}$.
Bárány, Fodor, Montejano, Oliveros, and Pór [BFM+14] proved a colorful $(p,q)$-theorem for families of convex sets in $\mathbb{R}^{d}$.
Theorem 1.12 (Bárány-Fodor-Montejano-Oliveros-Pór [BFM+14]: Colorful $(p,q)$-theorem).
For any three natural numbers $p\geq q\geq d+1$, $\exists M=M(p,q,d)$ such that the following holds: if $\mathcal{F}_{1},\dots,\mathcal{F}_{p}$ be finite families of convex sets in $\mathbb{R}^{d}$ satisfying colorful $(p,q)$-property with respect to points then there exists at least $q-d$ indices $i\in\left\{1,\dots,p\right\}$ such that $\mathcal{F}_{i}$ has a point transversal of size at most $M$.
Our results
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of compact convex sets in $\mathbb{R}^{d}$. An infinite sequence of compact convex sets $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a heterochromatic sequence of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ if there exists a strictly increasing sequence of natural numbers $\left\{i_{n}\right\}_{n\in\mathbb{N}}$ such that for all $n\in\mathbb{N}$ we have $B_{n}\in\mathcal{F}_{i_{n}}$. A heterochromatic sequence is said to be a strictly heterochromatic sequence if each family has a representative in the sequence, that is, a heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ is said to be a strictly heterochromatic sequence if for each $n\in\mathbb{N},\;B_{n}\in\mathcal{F}_{n}$.
A sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets in $\mathbb{R}^{d}$ is said to be $k$-dependent if there exists a $k$-flat $K$ and $k+2$ distinct integers $i_{1},\dots,i_{k+2}$ such that for all $j\in[k+2]$ we have $K\cap B_{i_{j}{}}\neq\emptyset.$ If a sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of convex sets is not $k$-dependent then we will call it $k$-independent. A collection $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ of families of convex sets in $\mathbb{R}^{d}$ is said to satisfy the heterochromatic $(\aleph_{0},k+2)$-property if every heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent. Similarly, we will say that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},k+2)$-property if every strictly heterochromatic sequence from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
We will prove a colorful variant of the Theorem 1.9 for families of closed convex sets satisfying heterochromatic $(\aleph_{0},k+2)$-property with respect to $k$-flats. To state the result we will need some additional notations and definitions. For a set $S\subseteq\mathbb{R}^{d}$, the out-radius $R_{S}$ is defined as the radius of the smallest closed ball containing $S$, and in-radius $r_{S}$ of $S$ is defined as the radius of the largest closed ball contained in $C$. The condition number $\sigma\left(S\right)$ of a set $S$ is defined as $\sigma(S):=\frac{r_{S}}{R_{S}}$. For a family $\mathcal{F}$ of subsets of $\mathbb{R}^{d}$, condition number $\sigma(\mathcal{F})$ of $\mathcal{F}$ is defined as the number $\sigma(\mathcal{F}):=\min\left\{\sigma(S)\;:\;S\in\mathcal{F}\right\}$.
Theorem 1.13 (Countably Colorful $(\aleph_{0},k+2)$-Theorem).
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed convex sets in $\mathbb{R}^{d}$ satisfying the heterochromatic $(\aleph_{0},k+2)$-property and
$\liminf\limits_{n\to\infty}\sigma(\mathcal{F}_{n})>0.$
Then there is a finite set of $k$-flats that pierces all but finitely many $\mathcal{F}_{n}$’s.
As a direct corollary of the above result we get the following generalization of Theorem 1.9 by Keller and Perles.
Corollary 1.14 (Generalization of Theorem 1.7 and Theorem 1.9).
Let $d$ be a natural number and $k$ be a non-negative integer $k$ with $10\leq k\leq d-$ 1. If $\mathcal{F}$ be a family of closed convex sets in $\mathbb{R}^{d}$ with $\sigma(\mathcal{F})>0$ and $\mathcal{F}$ satisfy $(\aleph_{0},k+2)$-property with respect to $k$-flats then $\mathcal{F}$ has a constant size $k$-transversal.
Using a different argument, Jung and Pálvölgyi [JP23] have given an alternate proof of the above corollary. Note that our proof uses entirely geometric arguments and builds on the approach of Keller and Perles [KP22]. We can also show that the convexity assumption in Theorem 1.13 and Corollary 1.14 cannot be weakened.
Theorem 1.15.
There is an infinite sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of compact connected sets in $\mathbb{R}^{2}$ with condition number $\rho>0$ such that $\{B_{n}\}_{n\in\mathbb{N}}$ satisfies the $(\aleph_{0},2)$ property, but $\{B_{n}\}_{n\in\mathbb{N}}$ is not pierceable by a finite number of points.
Observe that Theorem 1.7, Theorem 1.9, and Theorem 1.13 all require the convex bodies to have nice shape. In Theorem 1.7 and Theorem 1.9, we need the convex bodies to be $(r,R)$-fat and closed balls respectively, and in Theorem 1.13 we require the condition number of the families to be bounded away from zero. It is natural to ask if one can prove a result, similar to the above mentioned results, for some natural families of convex bodies that are neither fat nor do they have a lower bound on the condition number. Chakraborty, Ghosh and Nandi [CGN23a] have proved an $(\aleph_{0},k+2)$-theorem for axis-parallel boxes and axis-parallel $k$-flats. Axis-parallel boxes are neither fat nor do they have a lower bound on the condition number away from zero. The following impossibility result shows that
  • Such a result is not true if the boxes are not axis-parallel.
  • Assuming that the convex bodies have non-zero volumes is not sufficient to establish a $(\aleph_{0},\cdot)$-theorem.
  • A natural question related to Theorem 1.7 was, whether an $(\aleph_{0},k+2)$-theorem can be obtained for families of convex sets without the fatness assumption but the following impossibility result shows that this is not possible, even for $k=0$.
Theorem 1.16 (Impossibility result).
There exists an infinite family $\mathcal{F}$ of pairwise intersecting rectangles in $\mathbb{R}^{2}$ with bounded diameters and non-zero volumes that cannot be pierced by any finite collection of points in $\mathbb{R}^{2}$.
Remark 1.17.
Note that the proof of Theorem 1.16 can be modified for right-angled triangles in $\mathbb{R}^{2}$ with bounded diameter and non-zero volumes, see Theorem 5.1.
We say that a collection $\mathcal{F}$ of compact convex sets has bounded diameter if there exists an $R>0$ with $\text{diam}(C)<R$ for all $C\in\mathcal{F}$. We say a collection $\mathcal{F}$ of compact convex sets is unbounded if $\bigcup_{C\in\mathcal{F}}C$ is unbounded. As defined in [AGP02], we say $y\in\mathbb{S}^{d-1}$ to be a limiting direction of an unbounded collection $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{d}$ with bounded diameter if there exists an infinite sequence of points $\left\{s_{n}\right\}_{n\in\mathbb{N}}$ in $\mathbb{R}^{d}$ and an infinite sequence of sets $\left\{S_{n}\right\}_{n\in\mathbb{N}}$ from $\mathcal{F}$ satisfying the following conditions:
  • for all $n\in\mathbb{N}$, $S_{n}\in\mathcal{F}$,
  • for all $n\neq m$ in $\mathbb{N}$, $S_{n}\neq S_{m}$,
  • for all $n\in\mathbb{N}$, $s_{n}\in S_{n}$, and
  • $\lim\limits_{n\to\infty}\|s_{n}\|=\infty$ and $\lim\limits_{n\to\infty}\frac{s_{n}}{\|s_{n}\|}=y$
For an unbounded collection $\mathcal{F}$ of compact convex sets with bounded diameter in $\mathbb{R}^{d}$, limiting direction set $LDS(\mathcal{F})$ of $\mathcal{F}$ is defined as
$LDS(\mathcal{F}):=\left\{y\in\mathbb{S}^{d-1}\,:\,y\;\mbox{is a limiting direction of}\;\mathcal{F}\right\}.$
An unbounded collection $\mathcal{F}$ of compact convex sets with bounded diameter is called $k$-unbounded if the vector space spanned by set $LDS(\mathcal{F})$ has dimension at least $k$. As already mentioned earlier in the introduction, Santaló [San40] established the impossibility of getting a Helly Theorem for transversals of arbitrary convex sets with $k$-flats. Aronov, Goodman and Pollack [AGP02] proved the Helly Theorem for $k$-flats for $k$-unbounded families of arbitrary convex compact sets. We show the impossibility of getting $(\aleph_{0},k+2)$-Theorem for $k$-unbounded families of arbitrary convex compact bodies.
Theorem 1.18 (No-go theorem for $k$-unbounded families).
There exists an infinite family $\mathcal{F}$ of compact convex sets in $\mathbb{R}^{3}$ satisfying the following:
  • (a)
    $\mathcal{F}$ is one-unbounded,
  • (b)
    $\mathcal{F}$ satisfies the $(\aleph_{0},3)$-property with respect to lines, and
  • (c)
    $\mathcal{F}$ does not have a finite line transversal.
Given Theorem 1.13 it is natural to ask the following question:
Let $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ be an infinite sequence of families of nicely shaped closed convex sets satisfying
weak heterochromatic $(\aleph_{0},k+2)$-property for $10\leq k\leq d-$ 1. Does there exist a family $\mathcal{F}_{m}$
from the sequence $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ that has a finite size $k$-transversal?
We show the following colorful $(\aleph_{0},2)$-Theorem for families of $(R,r)$-fat convex sets satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Theorem 1.19 (Strongly Colorful $(\aleph_{0},2)$-Theorem).
Let $d\in\mathbb{N}$, and for each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of $(r,R)$-fat closed convex sets in $\mathbb{R}^{d}$. If $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property then there exist at least two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ with $i\neq j$ and each of them can be pierced by finitely many points.
The following proposition shows that Theorem 1.19 is tight for families satisfying weak heterochromatic $(\aleph_{0},2)$-property.
Proposition 1.20 (Tightness of Strongly Colorful Theorem 1.19).
For each $n\in\mathbb{N}$, there exists a family $\mathcal{F}_{n}$ of units balls in $\mathbb{R}^{2}$ such that $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ satisfies the following
  • (i)
    $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies the weak heterochromatic $(\aleph_{0},2)$-property, and
  • (ii)
    exactly two distinct families $\mathcal{F}_{i}$ and $\mathcal{F}_{j}$ from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ are pierceable by finite number of points, and the rest of the families from $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ are not pierceable by a finite number of points.
We show the impossibility of getting an analog of Theorem 1.19 for all natural numbers $k$ satisfying $0<k<d$ even if we restrict ourselves to families of unit balls.
Theorem 1.21 (No-go result for strongly colorful $(\aleph_{0},k+2)$-Theorem).
Suppose $d\in\mathbb{N}$ and $d>1$. Then for any non-negative integer $k$ with $11\leq k\leq d-$ 1, there exists a collection $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfying the following:
  • (i)
    for all $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ cannot be pierced by finitely many $k$-flats and
  • (ii)
    every strictly heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of $\left\{\mathcal{F}_{n}\right\}_{n\in\mathbb{N}}$ is $k$-dependent.
Organisation of the paper
  • In Section 2 we will give the definitions and notations used in this paper.
  • In Section 3 we will prove of Theorem 1.13.
  • In Section 4, we prove Theorem 1.19 and Proposition 1.20.
  • We will prove Theorem 1.15, Theorem 1.16, Theorem 1.18 and Theorem 1.21 in Section 5.
2 Definitions and notations
We will use the following definitions and notations in this paper.
  • For any $n\in\mathbb{N}$, $[n]$ denotes the set $\left\{1,\dots,n\right\}$.
  • For any $\alpha,\beta\in\mathbb{R}$ with $\alpha\leq\beta$, $[\alpha,\beta]$ denotes the interval $[\alpha,\beta]:=\left\{x\in\mathbb{R}:\;\alpha\leq x\leq\beta\right\}$.
  • Any axis-parallel box $B$ in $\mathbb{R}^{d}$ is a closed convex set of the form $B=[\alpha_{1},\beta_{1}]\times\dots\times[\alpha_{d},\beta_{d}]$, where $\alpha_{i},\beta_{i}\in\mathbb{R}$ with $\alpha_{i}\leq\beta_{i}$, $\forall i\in[d]$.
  • Let $0<r\leq R$. A subset $A$ of $\mathbb{R}^{d}$ is called $(r,R)$-fat if $A$ contains a closed ball of radius $r$ and is contained in a closed ball of radius $R$.
  • We will denote by $\mathcal{O}$ the origin in $\mathbb{R}^{d}$.
  • $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$ with center at the origin $\mathcal{O}$.
  • For any $x\in\mathbb{R}^{d}\setminus\{\mathcal{O}\}$, $\hat{x}$ denotes the central projection of $x$ on $\mathbb{S}^{d-1}$, that is, $\hat{x}:=\frac{x}{||x||}$.
  • For any $x\in\mathbb{R}^{d}$ and a ball $B$, $\text{dist}(x,B):=\min\{\text{dist}(x,b)\;|\;b\in B\}$.
  • For any $k$ such that $10\leq k\leq d-$ 1, any $k$-dimensional affine subspace of $\mathbb{R}^{d}$ is called a $k$-flat in $\mathbb{R}^{d}$. For example any point in $\mathbb{R}^{d}$ can be considered as a 00-flat, any line in $\mathbb{R}^{d}$ as a $1$-flat, any hyperplane in $\mathbb{R}^{d}$ as $(d-1)$-flat.
  • Direction of a $k$-flat $\mathcal{K}$ is defined as follows: First $\mathcal{K}$ is translated $\mathcal{K}^{\prime}$ so that $\mathcal{K}^{\prime}$ passes through the origin. Then direction of $\mathcal{K}$ is defined as the $(k-1)$-sphere at which $\mathcal{K}^{\prime}$ intersects $\mathbb{S}^{d-1}$. This definition has been taken from Aronov, Goodman and Pollack [AGP02].
  • A $k_{1}$-flat $\mathcal{K}_{1}$ and a $k_{2}$-flat $\mathcal{K}_{2}$ are said to be parallel if one of their directions contains the other, that is, if we translate $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ to $\mathcal{K}^{\prime}_{1}$ and $\mathcal{K}^{\prime}_{2}$ respectively so that both $\mathcal{K}^{\prime}_{1}$ and $\mathcal{K}^{\prime}_{2}$ pass through the origin then either $\mathcal{K}^{\prime}_{1}\subseteq\mathcal{K}^{\prime}_{2}$ or $\mathcal{K}^{\prime}_{2}\subseteq\mathcal{K}^{\prime}_{1}$. This definition has been taken from Keller and Perles [KP22].
  • For any $x\in\mathbb{R}^{d}$ and $\varepsilon>0$, $B(x,\varepsilon)$ denotes the closed ball of radius $\varepsilon$ with centered at the origin and $B^{o}(x,\varepsilon)$ denotes the open ball of radius $\varepsilon$ with centered at the origin. $B^{o}(x,\varepsilon)$ is also considered as the (open) $\varepsilon$-neighbourhood of $x$.
  • For any $x\in\mathbb{S}^{d-1}$ and $\varepsilon>0$, an $\varepsilon$-neighbourhood of $x$ on the sphere $\mathbb{S}^{d-1}$ is $B^{o}(x,\varepsilon)\cap\mathbb{S}^{d-1}$.
  • For any $\hat{x}\in\mathbb{S}^{d-1}$ and $\varepsilon>0$ and any family of closed balls $\mathcal{F}$ in $\mathbb{R}^{d}$ the (open) $\varepsilon$-neighbourhood of $\hat{x}$ in $\mathcal{F}$ is $\mathcal{F}_{(\hat{x},\varepsilon)}=\{B(y,r)\in\mathcal{F}\;|\;\text{dist}(\hat{y},\hat{x})<\varepsilon\}$.
  • An unbounded cone $\mathcal{C}$ in $\mathbb{R}^{d}$ with vertex at the origin $\mathcal{O}$ is a convex set such that for any $u,v\in\mathcal{C}$ and $\alpha,\beta\geq 0$, we have $\alpha u+\beta v\in\mathcal{C}$. The axis of $\mathcal{C}$ is a straight line $\ell$ passing through the vertex such that $\mathcal{C}$ has a circular symmetry about $\ell$. If we translate $\mathcal{C}$ at a point $z\in\ell$, we get the unbounded cone $\mathcal{C}^{\prime}=\{z+c\;|\;c\in\mathcal{C}\}$.
  • Limit inferior of a sequence $\left\{x_{n}\right\}_{n\in\mathbb{N}}$ of real numbers is defined as
    $\liminf_{n\to\infty}x_{n}:=\sup\left\{\inf\left\{x_{m}\;:\;m\geq n\right\}\;:\;n\in\mathbb{N}\right\}.$
  • A collection of $k$-flats $\mathcal{K}$ is said to be an $\varepsilon$-net cover of a region $P\subsetneq\mathbb{R}^{d}$ if for each point $p\in P$, there exists a $k$-flat $\kappa\in\mathcal{K}$ such that $\text{dist}(p,\kappa)<\varepsilon$.
  • $k$-set: We say $\mathcal{F}$ is a $k$-set if $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ which does not have a finite $k$-flat transversal. We say $A\subseteq\mathbb{R}^{d}$ contains a $k$-set $\mathcal{B}$ or $\mathcal{B}$ is contained in $A$ if $\forall B\in\mathcal{B}$, $B\subseteq A$.
[…]
  • For any $S\in\mathcal{F}$ where $\mathcal{F}$ is a collection of sets in $\mathbb{R}^{d}$, $B_{S}$ denotes a largest ball inscribed into $S$ and $x_{S}$ denotes the centre of $B_{S}$.
  • For any $S\in\mathcal{F}$ where $\mathcal{F}$ is a collection of compact sets in $\mathbb{R}^{d}$, $r_{S}$ and $R_{S}$ denote the radius of the largest closed ball contained in $S$ and the radius of the smallest closed ball containing $S$ respectively.
  • For a family $\mathcal{F}$ of sets in $\mathbb{R}^{d}$ and any two positive numbers $r,R\in\mathbb{R}^{+}$, $\mathcal{F}_{r,R}:=\{S\in\mathcal{F}\mid r_{S}\geq r\text{ and }R_{S}\leq R\}$, $\mathcal{F}_{<r}:=\{S\in\mathcal{F}\mid r_{S}<r\}$ and $\mathcal{F}_{>R}:=\{S\in\mathcal{F}\mid R_{S}>R\}$.
3 Convex sets with bounded condition number
In this section we prove Theorem 1.13 in the following three steps:
  • first we show that heterochromatic $(\aleph_{0},.)$-theorem implies the existence of a finite hitting set,
  • next we establish that monochromatic $(\aleph_{0},k+2)$-theorem for compact convex bodies and $k$-flats implies Theorem 1.13, and
  • finally we prove the monochromatic version of Theorem 1.13.
3.1 Heterochromatic $(\aleph_{0},.)$-theorem implies the existence of a finite hitting set
In this section, we shall show that the heterochromatic $(\aleph_{0},q)$-theorem with $q\in\mathbb{N}$ implies the existence of a finite hitting set, regardless of the types of sets chosen from $\mathbb{R}^{d}$. Let us first prove a Lemma. Let $\mathcal{B}$ and $\mathcal{S}$ be collections of nonempty sets in $\mathbb{R}^{d}$. For any $\mathcal{F}\subseteq\mathcal{B}$, we say $\mathcal{F}$ is finitely pierceable by $\mathcal{S}$ if there exists a finite set $S\subseteq\mathcal{S}$ such that for all $B\in\mathcal{F}$, we have $\bigcup_{A\in S}A\cap B\neq\emptyset$. A sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ $\forall n\in\mathbb{N}$ is said to have the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$ if for every heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ chosen from $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, we have a set $A\in\mathcal{S}$ such that there are $q$ distinct $B_{n}$’s for which $B_{n}\cap A\neq\emptyset$. We say that the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, there exists an $n\in\mathbb{N}$ for which $\mathcal{F}_{n}$ is finitely pierceable by $\mathcal{S}$. We say that the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, all but finitely many $\mathcal{F}_{n}$’s are finitely pierceable by $\mathcal{S}$.
Lemma 3.1 (Heterochromatic Property Lemma).
The heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ implies the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$.
Proof.
Let the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ hold. Let $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, $\mathcal{F}_{n}\subseteq\mathcal{B}$, be a sequence that satisfies the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$. If possible, let $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ where $\{n_{i}\}_{i\in\mathbb{N}}$ is a strictly increasing sequence in $\mathbb{N}$ be such that no $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. As $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ also satisfies the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, there is an $i\in\mathbb{N}$ for which $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. But this is a contradiction, and therefore, the result follows. ∎
So, we shall be using the phrases heterochromatic $(\aleph_{0},q)$-theorem and infinite heterochromatic $(\aleph_{0},q)$-theorem interchangeably. It is trivially seen that for some $\mathcal{B}$ and $\mathcal{S}$ satisfying the heterochromatic $(\aleph_{0},q)$-theorem and any $N\in\mathbb{N}$, we can find a sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ that satisfy the heterochromatic $(\aleph_{0},q)$ property and for $N$ distinct values of $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is not finitely pierceable by $\mathcal{S}$. A simple example is found by setting $\mathcal{B}$ to be the set of all unit balls in $\mathbb{R}^{d}$, $\mathcal{S}$ to be the set of all points in $\mathbb{R}^{d}$ and $\mathcal{F}_{1}=\dots=\mathcal{F}_{N}=\mathcal{B}$ and $\mathcal{F}_{n}=\{B(\mathcal{O},1)\}$ for all $n\in\mathbb{N}\setminus\{1,\dots,N\}$.
We can show that the heterochromatic $(\aleph_{0},k+2)$-theorem implies the existence of a finite hitting set.
Lemma 3.2 (Heterochromatic Theorem and a Finite Hitting Set).
Let $\mathcal{B}$ and $\mathcal{S}$ be collections of sets in $\mathbb{R}^{d}$ and $q\in\mathbb{N}$, where $\mathcal{B}$ satisfies the heterochromatic $(\aleph_{0},q)$-theorem with respect to $\mathcal{S}$. If $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ be a sequence of subsets of $\mathcal{B}$ that satisfies the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, then there exists an $N\in\mathbb{N}$ and a finite collection $\mathcal{K}\subset\mathcal{S}$ such that $\forall n>N$ where $n\in\mathbb{N}$, $\forall C\in\mathcal{F}_{n}$, we have $\bigcup_{K\in\mathcal{K}}K\cap C\neq\emptyset$.
Proof.
By Lemma 3.1 we have that there is an $N\in\mathbb{N}$ such that every $\mathcal{F}_{n}$ with $n>N$ is finitely pierceable by $\mathcal{S}$. Let $\mathcal{F}=\bigcup_{n=N}^{\infty}\mathcal{F}_{n}$. Then for any infinite set $F\subseteq\mathcal{F}$, either $F$ contains infinitely many elements of some $\mathcal{F}_{n}$ with $n\geq N$, or $F$ contains a heterochromatic sequence with respect to the sequence $\{\mathcal{F}_{n}\}_{n=N}^{\infty}$. In either case, this means there is an $S\in\mathcal{S}$ that pierces $q$ elements of $F$. Since the heterochromatic $(\aleph_{0},q)$-theorem implies the monochromatic $(\aleph_{0},q)$-theorem, $\mathcal{F}$ must be finitely pierceable by $\mathcal{S}$. ∎
The above is a generalization of the arguments given in [CGN23b].
3.2 Monochromatic $(\aleph_{0},k+2)$-theorem implies the heterochromatic version
Here we show that the monochromatic $(\aleph_{0},k+2)$-theorem for compact convex bodies and $k$-flats implies the heterochromatic $(\aleph_{0},k+2)$-theorem for the same. In order to do that, we shall need the following lemma by Keller and Perles.
Lemma 3.3 ([KP22]).
Let $\mathcal{F}$ be a family of compact convex sets in $\mathbb{R}^{d}$, $10\leq k\leq d-$ 1 and $m\in\mathbb{N}$. If every finite subset of $\mathcal{F}$ can be pierced by $m$ $k$-flats, then $\mathcal{F}$ can be pierced by $m$ $k$-flats.
Let $\mathcal{B}$ be a collection of compact convex sets in $\mathbb{R}^{d}$, and $0\leq k<d$. We say $\mathcal{B}$ satisfies the $(\aleph_{0},k+2)$-theorem with respect to $k$-flats if for all $\mathcal{F}\subseteq\mathcal{B}$, $\mathcal{F}$ satisfying the $(\aleph_{0},k+2)$-property implies that $\mathcal{F}$ is pierced by finitely many $k$-flats, then we say that the $(\aleph_{0},k+2)$-theorem holds for $\mathcal{B}$ with respect to the set of all $k$-flats.
Lemma 3.4.
Let $\mathcal{B}$ be a collection of compact convex sets in $\mathbb{R}^{d}$ and $0\leq k<d$. Then if the $(\aleph_{0},k+2)$-theorem holds for $\mathcal{B}$ with respect to the set of all $k$-flats, then the heterochromatic $(\aleph_{0},k+2)$-theorem also holds for $\mathcal{B}$ with respect to the set of all $k$-flats.
Proof.
Let $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ be a sequence of families of sets in $\mathcal{B}$ and for every $n\in\mathbb{N}$, $\mathcal{F}_{n}\subseteq\mathcal{B}$ be not pierceable by finitely many $k$-flats. Then by Lemma 3.3, we can obtain, for every $n\in\mathbb{N}$, a finite set $S_{n}\subset\mathcal{F}_{n}$ such that $S_{n}$ is not pierceable by $n$ $k$-flats. Let $S=\bigcup_{n\in\mathbb{N}}S_{n}$. Then $S$ is not pierceable by finitely many $k$-flats. Since $S\subset\mathcal{B}$ and $\mathcal{B}$ satisfies the $(\aleph_{0},k+2)$-theorem with respect to $k$-flats, we have an infinite $k$-independent sequence $\{B_{n}\}_{n\in\mathbb{N}}$ in $S$. As one can find a heterochromatic subsequence $\{B_{n_{i}}\}_{i\in\mathbb{N}}$ such that $B_{n_{i}}\in\mathcal{F}_{n_{i}}$ for all $i\in\mathbb{N}$ for some strictly increasing sequence $\{i_{n}\}_{n\in\mathbb{N}}$ in $\mathbb{N}$, we see that $\mathcal{B}$ satisfies the heterochromatic $(\aleph_{0},k+2)$-theorem with respect to the set of all $k$-flats. ∎
3.3 Monochromatic version of Theorem 1.13
Now we shall prove the monochromatic version of Theorem 1.13. As is common in several proofs of such results concerning higher dimensional flats (for example, see [AGP02, KP22]), we prove the result using induction.
Theorem 3.5 ($(\aleph_{0},k+2)$-theorem for sets with bounded condition number).
Suppose $\mathcal{F}$ is a family of compact convex sets in $\mathbb{R}^{d}$ with condition number $\sigma(F)\geq\sigma,\;\forall F\in\mathcal{F}$. Then either $\mathcal{F}$ is pierceable by a finite number of $k$-flats or $\mathcal{F}$ contains a $k$-independent sequence.
Observation 3.6.
Given a $k$-set $\mathcal{F}$ with condition number $\sigma>0$, we can safely assume that for any $r,R>0$ with $r<R$, $\mathcal{F}_{r,R}=\{S\in\mathcal{F}\;|\;r_{S}\geq r\;\&\;R_{S}\leq R\}$ is not a $k$-set. For otherwise, using Theorem 1.7 we can find a $k$-independent sequence from this $(r,R)$-fat subfamily.
So from now on, we assume that $\mathcal{F}_{r,R}$ is not a $k$-set, for any $r,R>0$ with $r<R$.
Observation 3.7.
For any $k$-set $\mathcal{F}$ with condition number $\sigma>0$, at least one of the following must be true:
  • (1)
    For every $r>0$, $\mathcal{F}_{<r}:=\{S\in\mathcal{F}\;|\;r_{S}<r\}$ is a $k$-set.
  • (2)
    For every $R>0$, $\mathcal{F}_{>R}:=\{S\in\mathcal{F}\;|\;R_{S}>R\}$ is a $k$-set.
Proof.
For any $r,R>0$ with $r<R$, we have $\mathcal{F}=\mathcal{F}_{<r}\cup\mathcal{F}_{r,R}\cup\mathcal{F}_{>R}$. Since we have already assumed that, $\mathcal{F}_{r,R}$ is not a $k$-set, so for a $k$-set $\mathcal{F}$, at least one of $\mathcal{F}_{<r}$ or $\mathcal{F}_{>R}$ must be a $k$-set. Now if for some $r>0$, $\mathcal{F}_{<r}$ is not a $k$-set, then $\mathcal{F}_{>R}$ must be a $k$-set for every $R>r$, and $\mathcal{F}_{>R}\subseteq\mathcal{F}_{>r^{\prime}}$ for every $0<r^{\prime}\leq r<R$. So in this case, $\mathcal{F}_{>R}$ is a $k$-set for all $R>0$. The other case is proven similarly. ∎
Observation 3.8.
Let $\mathcal{F}$ be a $k$-set with condition number $\sigma>0$ such that there is a bounded convex open set $A$ and $\overline{A}$ intersects every set in $\mathcal{F}$. Then for every $R>0$, $\mathcal{F}_{>R}$ is not a 00-set.
Proof.
Here we show that for every $R>0$, $\mathcal{F}_{>R}$ is pierceable by finitely many points. Recall from the definitions that for every $S\in\mathcal{F}_{>R}$, $B_{S}$ denotes a largest inscribed ball in $S$. It is easy to see that there are a finite number of points $P$ such that for every $S\in\mathcal{F}_{>R}$ with $B_{S}\cap\overline{A}\neq\emptyset$, we have $S\cap P$. That is because the radius of $B_{S}$ is at least $\frac{R}{\sigma}$, and all these $B_{S}$ intersect $\overline{A}$. Now, let $\mathcal{S}\subseteq\mathcal{F}_{>R}$ be the collection of all sets $S$ in $\mathcal{F}_{>R}$ for which $B_{S}$ does not intersect $\overline{A}$. Then there is an $\varepsilon>0$ such that for every $S\in\mathcal{S}$, there is a point $p_{S}$ on $\partial A\cap S$ from which a finite cone $C_{S}$ of aperture $\varepsilon$ and axis length $\frac{R}{\sigma}$ exists that lies completely within $S$. It is easily seen that we can find a finite set $Q\subsetneq\mathbb{R}^{d}$ such that for all $S\in\mathcal{S}$, $Q\cap C_{S}\neq\emptyset$, since for any two $S_{1},S_{2}\in\mathcal{S}$, $C_{S_{1}}$ and $C_{S_{2}}$ are identical up to translation and rotation, every $C_{S}$ for $S\in\mathcal{S}$ contains a ball with nonzero volume and there is a bounded region that contains $C_{S}$ for every $S\in\mathcal{S}$. ∎
Claim 3.9.
Given a $k$-set $\mathcal{F}$ with condition number $\sigma>0$, there is a point $p\in\mathbb{R}^{d}$ and a ray $\vec{r}$ from $p$ such that one of the following two cases must arise:
  • (1)
    $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in R(\vec{r},\frac{1}{n})\}$ is a $k$-set.
  • (2)
    $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in Q(\vec{r},n)\}$ is a $k$-set.
Proof.
Suppose $\mathcal{G}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in\mathbb{R}^{d}\setminus B(\mathcal{O},n)\}$. Then either $\forall n\in\mathbb{N},\;\mathcal{G}_{n}$ is a $k$-set or $\exists n^{\prime}\in\mathbb{N}$ such that $\mathcal{G}_{n^{\prime}}$ is not a $k$-set.
Case I: $\forall n\in\mathbb{N},\;\mathcal{G}_{n}$ is a $k$-set.
[…]
For $\mathcal{G}_{n}$, we set
[…]
Clearly, for each $n\in\mathbb{N}$, $\bigcap_{i=1}^{n}V_{i}\neq\emptyset$. Since each $V_{i}$ is compact, this implies that $\bigcap_{n\in\mathbb{N}}V_{n}\neq\emptyset$. Now choose $v\in\bigcap_{n\in\mathbb{N}}V_{n}$. Let $\vec{r}:=\{cv\;|\;c\geq 0\}$. Then we get that, for each $n\in\mathbb{N}$, $\mathcal{F}_{n}=\left\{S\in\mathcal{F}\;|\;x_{S}\in Q(\vec{r},n)\right\}$ is a $k$-set.
Case II: $\exists n^{\prime}\in\mathbb{N}$ such that $\mathcal{G}_{n^{\prime}}$ is not a $k$-set.
We claim that, $\exists p\in\bar{B}(\mathcal{O},n^{\prime})$ such that for every open ball $B$ with $p\in B$, we have $\mathcal{F}_{B}=\{S\in\mathcal{F}\;|\;x_{S}\in B\}$ is a $k$-set. Otherwise, for every $a\in\bar{B}(\mathcal{O},n^{\prime})$, we have an open ball $B_{a}$ with $a\in B_{a}$ such that $\mathcal{F}_{B_{a}}$ is not a $k$-set. Again by compactness of $\bar{B}(\mathcal{O},n^{\prime})$ we get that the open cover $\{B_{a}\;|\;a\in\bar{B}(\mathcal{O},n^{\prime})\}$ of $\bar{B}(\mathcal{O},n^{\prime})$ has a finite subcover $B_{a_{1}},\dots,B_{a_{t}}$. This implies that $\mathcal{G}^{\prime}_{n^{\prime}}=\{S\in\mathcal{F}\;|\;x_{S}\in B(\mathcal{O},n^{\prime})\subset\cup_{i=1}^{t}B_{a_{i}}\}$ is not a $k$-set. Again by our assumption $\mathcal{G}_{n^{\prime}}$ is not a $k$-set and hence so is $\mathcal{F}=\mathcal{G}_{n^{\prime}}\cup\mathcal{G}^{\prime}_{n^{\prime}}$. But this contradicts the fact that $\mathcal{F}$ is a $k$-set.
[…]
Claim 3.10.
Let $\mathcal{F}$ be a $k$-set with condition number $\sigma>0$ where $10<k\leq d-$ 1. Suppose $p\in\mathbb{R}^{d}$ and $\vec{r}$ is a ray from $p$ such that $\forall n\in\mathbb{N}$, $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ is a nested sequence of $k$-sets with $\mathcal{F}_{n}\subseteq\mathcal{F}$, and for all $S\in\mathcal{F}_{n}$, we have $x_{S}\in R(\vec{r},\frac{1}{n})$. Then $\mathcal{G}_{n}:=\{S\in\mathcal{F}_{n}\;|\;S\subset R(\vec{r},\frac{1}{n})\}$ is a $k$-set with condition number $\sigma>0$.
Proof.
From Observation 3.8, we know that for every $R>0$, $\mathcal{F}_{>R}\cap\mathcal{F}_{n}$ is not a $k$-set. Since we have assumed that for every $r,R>0$ with $r<R$, $\mathcal{F}_{r,R}$ is not a $k$-set, $\forall r>0,\;\mathcal{F}_{n}\cap\mathcal{F}_{<r}$ must be a $k$-set for every $r>0$. Without loss of generality, we assume that $p$ is the origin $\mathcal{O}$ and $\vec{r}$ is the positive $d$-th axis, i.e, $\vec{r}=\{c(0,\dots,0,1)\in\mathbb{R}^{d}\;|\;c\geq 0\}$. We also assume that $\forall S\in\mathcal{F},\;\vec{r}\cap S=\emptyset$. Now we define a function $m:\mathbb{N}\rightarrow\mathbb{N}$ by $m(n)=\left\lceil 2\sigma(2^{n}+2\sigma)\right\rceil$ and pick any $n\in\mathbb{N}$. We choose any $S\in\mathcal{F}_{m(n)}$. Then $r_{S}<\text{dist}(x_{S},\vec{r})$. Now we take the $u\in\vec{r}$ for which $\text{dist}(x_{S},\vec{r})=\text{dist}(x_{S},u)$. Clearly, $\frac{||x_{S}-u||}{||u||}\leq\frac{1}{m(n)}$. Then
$\frac{r_{S}}{||u||}\leq\frac{1}{m(n)}\implies\frac{2R_{S}}{||u||}\leq\frac{2\sigma}{m(n)}\leq\frac{1}{2^{n}}.$
Also,
$\frac{2R_{S}}{||u||-2R_{S}}\leq\frac{2\sigma}{m(n)-2\sigma}\leq\frac{1}{2^{n}}.$
This shows that $S\subset R(\vec{r},\frac{1}{n})$. So $\mathcal{G}_{n}\supseteq\mathcal{F}_{m(n)}$ and hence $\mathcal{G}_{n}$ is a $k$-set. ∎
Now we are ready to prove Theorem 3.5. First we prove the case when $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\left\{S\in\mathcal{F}\;|\;x_{S}\in R(\vec{r},\frac{1}{n})\right\}$ is a $k$-set.
Lemma 3.11.
Suppose $\mathcal{F}$ is a family of compact convex $k$-sets in $\mathbb{R}^{d}$ with condition number $\sigma(S)\geq\sigma,\;\forall S\in\mathcal{F}$ and there is a point $p\in\mathbb{R}^{d}$ and a ray $\vec{r}$ from $p$ such that $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in R(\vec{r},\frac{1}{n})\}$ is a $k$-set. Then $\mathcal{F}$ contains a $k$-independent sequence.
Proof.
We shall prove the result by induction on $k$ and $d$.
Base case: $k=0$ and $d\in\mathbb{N}$.
$\mathcal{F}$ is a family of compact convex 00-sets in $\mathbb{R}^{d}$ with condition number $\sigma(S)\geq\sigma\;\forall S\in\mathcal{F}$ and there is a point $p\in\mathbb{R}^{d}$ and a sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ of 00-sets such that $\forall n\in\mathbb{N},\mathcal{F}_{n+1}\subseteq\mathcal{F}_{n}\subseteq\mathcal{F}$ and $S\in\mathcal{F}_{n}$ implies $x_{S}\in B(p,\frac{1}{n})$. Then we shall show that, $\mathcal{F}$ contains a 00-independent sequence $\{S_{n}\}_{n\in\mathbb{N}}$ with $S_{n}\in\mathcal{F}_{n}$ and $p\notin S_{n}$ for each $n\in\mathbb{N}$. From Observation 3.8, we have that $\mathcal{F}_{2}\setminus\mathcal{F}_{>\frac{1}{2}}$ is also a 00-set. Pick a $S_{1}$ from $\mathcal{F}_{2}\setminus\mathcal{F}_{>\frac{1}{2}}$ such that $p$ does not intersect $S_{1}$. Now there is an $n_{1}\in\mathbb{N}$ such that $B(p,\frac{1}{n_{1}})$ does not intersect $S_{1}$. As $\mathcal{F}_{4n_{1}}\setminus\mathcal{F}_{>\frac{1}{4n_{1}}}$ is a 00-set, pick a $S_{2}$ from this set that does not contain $p$. Now, as any point in $S_{2}$ is within $\frac{1}{4n_{1}}+\frac{2}{4n_{1}}=\frac{3}{4n_{1}}$ distance from $p$, $S_{2}$ is contained in $B(p,\frac{1}{n_{1}})$. Again, there is an $n_{2}\in\mathbb{N}$ such that $B(p,\frac{1}{n_{2}})$ does not intersect $S_{2}$. Continuing this way, we obtain the sequence $\{S_{n}\}_{n\in\mathbb{N}}$ such that (1) is satisfied. Note that if we have a 00-set $\mathcal{F}$, then if it satisfies Claim 3.9(1), then Claim 3.9 gives us a sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfying the properties in Lemma 3.11. So, Lemma 3.11 is proven for $k=0$.
Induction hypothesis:
$\mathcal{F}$ is a family of compact convex $(k-1)$-sets in $\mathbb{R}^{d}$ with condition number $\sigma(S)\geq\sigma,\;\forall S\in\mathcal{F}$ and there is a point $p\in\mathbb{R}^{d}$ and a sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ of $(k-1)$-sets such that $\forall n\in\mathbb{N},\mathcal{F}_{n+1}\subseteq\mathcal{F}_{n}\subseteq\mathcal{F}$ and $S\in\mathcal{F}_{n}$ implies $x_{S}\in B(p,\frac{1}{n})$. Then $\mathcal{F}$ contains a $(k-1)$-independent sequence $\{S_{n}\}_{n\in\mathbb{N}}$ with $S_{n}\in\mathcal{F}_{n}$ for each $n\in\mathbb{N}$ satisfying $p\not\in\mathcal{C}_{k}(S_{i_{1}},\dots,S_{i_{k}})$, where $\mathcal{C}_{k}(S_{i_{1}},\dots,S_{i_{k}})$ is the union of all $(k-1)$-flats passing through $S_{i_{1}},\dots,S_{i_{k}}$.
Inductive step:
Now we induct on $k$ and $d$. We have already proved the result when $k=0$ in the proof of the base case. So let $k>0$. Suppose we have $p\in\mathbb{R}^{d}$ and a ray $\vec{r}$ from $p$ such that $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in R(\vec{r},\frac{1}{n})\}$ is a $k$-set. By Claim 3.10, we get that $\mathcal{G}_{n}:=\{S\in\mathcal{F}_{n}\;\mid S\subset R(\vec{r},\frac{1}{n})$ is a $k$-set. Then using Lemma 3.3 we obtain a countably infinite collection $\mathcal{G}\subseteq\mathcal{F}$ such that $\forall n\in\mathbb{N},\;\mathcal{G}\setminus\mathcal{F}_{n}$ is finite and $\mathcal{G}$ is a $k$-set. Without loss of generality, we assume that $p$ is the origin $\mathcal{O}$ and $\vec{r}$ is the positive $d$-th co-ordinate axis. Let $Y$ be the hyperplane perpendicular to $\vec{r}$ passing through $\mathcal{O}$. Then by the induction hypothesis there exists a sequence $\{S_{n}\}_{n\in\mathbb{N}}\subseteq\mathcal{G}$ such that the sequence of the projections of each $S_{n}$ on $Y$ is a $(k-1)$-independent sequence and no $(k-1)$-flat on $Y$ passing through the projections of $k$ distinct $S_{n}$’s contains the origin $\mathcal{O}$. Since for every $n\in\mathbb{N}$, $\mathcal{G}\setminus\mathcal{F}_{n}$ is finite, using Claim 3.10 on $\mathcal{G}_{n}:=\mathcal{G}\cap\mathcal{F}_{n}$ we can assume without loss of generality that for every $n\in\mathbb{N}$, $S_{n}$ is contained in $R(\vec{r},\frac{1}{n})$. Now we pick $k$ distinct sets $S_{i_{1}},\dots,S_{i_{k}}$ from $\{S_{n}\}_{n\in\mathbb{N}}$. Consider the union of all $k$-flats passing through the origin $\mathcal{O}$ and $S_{i_{1}},\dots,S_{i_{k}}$, denoted by $\mathcal{C}_{k+1}(S_{i_{1}},\dots,S_{i_{k}},\{\mathcal{O}\})$. Then we can show that $\mathcal{C}_{k+1}(S_{i_{1}},\dots,S_{i_{k}},\{\mathcal{O}\})\cap\vec{r}=\{\mathcal{O}\}$. That is because if there is a $k$-flat through $S_{i,1},\dots,S_{i_{k}}$ that contains $\vec{r}$, then its projection onto $Y$ is a $(k-1)$-flat that passes through $cO$, which is a contradiction. This means $\exists n_{k+1}\in\mathbb{N}$ such that $\mathcal{C}_{k+1}(S_{i_{1}},\dots,S_{i_{k}},\{\mathcal{O}\})\cap R(\vec{r},\frac{1}{n_{k+1}})=\{\mathcal{O}\}$. We pick $S_{i_{k+1}}\in\{S_{n}\}_{n\in\mathbb{N}}$ such that $i_{k+1}>i_{j}\;\forall j\in[k]$ and $S_{i_{k+1}}\subset R(\vec{r},\frac{1}{n_{k+1}})$. So $\mathcal{O}\not\in\mathcal{C}_{k+1}(S_{i_{1}},\dots,S_{i_{k}},S_{i_{k+1}})$.
Now let $S_{i_{1}},\dots,S_{i_{m}}$ be chosen from $\{S_{n}\}_{n\in\mathbb{N}}$ with $i_{1}<i_{2}<\dots<i_{m}$ such that for any distinct $k+1$ sets $S^{\prime}_{1},\dots,S^{\prime}_{k+1}$ chosen from $S_{i_{1}},\dots,S_{i_{m}}$ we have $\mathcal{O}\not\in\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})$ and for any $k+2$ distinct sets $S^{\prime}_{1},\dots,S^{\prime}_{k+2}$ chosen from $S_{i_{1}},\dots,S_{i_{m}}$ we have $\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})\cap S^{\prime}_{k+2}=\emptyset$. Since for any distinct $k+1$ sets $S^{\prime}_{1},\dots,S^{\prime}_{k+1}$, $\mathcal{O}\not\in\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})$, there exists $n_{1}\in\mathbb{N}$ such that $B(\mathcal{O},n_{1})\cap\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})=\emptyset$. Also for any distinct $k$ sets $S^{\prime}_{1},\dots,S^{\prime}_{k}$ chosen from $S_{i_{1}},\dots,S_{i_{m}}$, we have $\exists n_{2}\in\mathbb{N}$ such that $R(\vec{r},\frac{1}{n_{2}})\cap\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k},\{\mathcal{O}\})=\{\mathcal{O}\}$. Now we pick $S_{i_{m+1}}\in\{S_{n}\}_{n\in\mathbb{N}}$ such that $i_{m+1}>i_{m}$ and $S_{i_{m+1}}\subset R(\vec{r},\frac{1}{\bar{n}})\setminus\{\mathcal{O}\}$, where $\bar{n}=\max\{n_{1},n_{2}\}$. Continuing this way, we obtain our required $k$-independent sequence. ∎
Next we prove the case when $\forall n\in\mathbb{N},\;\mathcal{F}_{n}:=\{S\in\mathcal{F}\;|\;x_{S}\in Q(\vec{r},n)\}$ is a $k$-set. Before that we require the following technical lemma.
Claim 3.12.
Let $\mathcal{F}$ be a $k$-set with condition number $\sigma>0$ where $10<k\leq d-$ 1. Suppose $p\in\mathbb{R}^{d}$ and $\vec{r}$ is a ray from $p$ such that $\forall n\in\mathbb{N}$, $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ is a nested sequence of $k$-sets with $\mathcal{F}_{n}\subseteq\mathcal{F}$, and for all $S\in\mathcal{F}_{n}$, we have $x_{S}\in Q(\vec{r},n)$. Then $\forall n\in\mathbb{N},\;\mathcal{G}_{n}=\{S\in\mathcal{F}_{n}\;|\>S\subset Q(\vec{r},n)\}$ is a $k$-set.
Proof.
From Observation 3.7, either we have $\forall r>0,\;\mathcal{F}_{<r}$ is a $k$-set or $\forall R>0,\;\mathcal{F}_{>R}$ is a $k$-set. Observe that $\forall r>0$, $\mathcal{F}_{<r}$ is a $k$-set. So now we assume that $\forall R>0,\;\mathcal{F}_{>R}$ is a $k$-set. Without loss of generality, let $p$ be the origin $\mathcal{O}$ and $\vec{r}$ be the positive $d$-th axis. Also assume that $\forall S\in\mathcal{F}_{>R},\;S\cap\vec{r}=\emptyset$. Now we define $\forall n\in\mathbb{N},\;m(n)=\left\lceil 2\sigma(2^{n}+2\sigma)\right\rceil$ and choose any $S\in\mathcal{F}_{m(n)}$. Let $u\in\vec{r}$ be such that $\text{dist}(x_{S},\vec{r})=\text{dist}(x_{S},u)$. Since $S\cap\vec{r}=\emptyset$, we have $r_{S}\leq||x_{S}-u||$. Also
$\frac{||x_{S}-u||}{||u||}\leq\frac{1}{m(n)}\implies\frac{r_{S}}{||u||}\leq\frac{1}{m(n)}\implies\frac{2R_{S}}{||u||}\leq\frac{2\sigma}{m(n)}\leq\frac{1}{2^{n}}.$
Again, observe that
$\frac{2R_{S}}{||u||-2R_{S}}\leq\frac{1}{2^{n}}.$
This shows that $\mathcal{F}_{m(n)}\subseteq\mathcal{G}_{n}$ and hence $\mathcal{G}_{n}$ is a $k$-set. ∎
Now we prove the Theorem 3.5.
See 3.5
Proof.
Suppose $\mathcal{F}$ is a family of compact convex $k$-sets in $\mathbb{R}^{d}$ with condition number $\sigma(S)\geq\sigma,\;\forall S\in\mathcal{F}$. We will show that $\mathcal{F}$ contains a $k$-independent infinite sequence. We will use induction on $d$ and $k$.
Base case: $d\in\mathbb{N}$ and $k=0$.
We will now show that if $\mathcal{F}$ is a 00-set with bounded condition number $\sigma(S)\geq\sigma,\;\forall S\in\mathcal{F}$, then $\mathcal{F}$ contains an infinite sequence of pairwise disjoint sets.
From Claim 3.9, we know that $\mathcal{F}$ either satisfies Claim 3.9(1), or Claim 3.9(2). If $\mathcal{F}$ satisfies Claim 3.9(1), then from the proof of the base case of Lemma 3.11, we get an 00-independent sequence in $\mathcal{F}$. So let us assume that $\mathcal{F}$ does not satisfy Claim 3.9(1) and therefore it satisfies Claim 3.9(2). Choose any $S_{1}\in\mathcal{F}_{1}$, where for every $n\in\mathbb{N}$, $\mathcal{F}_{n}$ is as defined in Claim 3.9(1). There is an $n_{1}\in\mathbb{N}$ for which $S_{1}\subset B(\mathcal{O},n_{1})$. If every set in $\mathcal{F}_{n_{1}}$ intersects $B(\mathcal{O},n_{1})$, then by Observation 3.8, $\mathcal{F}_{n_{1}}\setminus\mathcal{F}_{<1}$ is a 00-set. But this means that for every $S\in\mathcal{F}_{n_{1}}$, $S\subset B(\mathcal{O},n_{1}+2)$, which contradicts our assumption that $\mathcal{F}$ satisfies Claim 3.9(2). Therefore, we can choose an $S_{2}\in\mathcal{F}_{n_{1}}$ such that $S_{2}\cap B(\mathcal{O},n_{1})=\emptyset$. Continuing this way, one can obtain a pairwise disjoint sequence $\{S_{n}\}_{n\in\mathbb{N}}$ with $S_{n}\in\mathcal{F}$.
Induction hypothesis:
If $\mathcal{F}$ is a $(k-1)$-set with condition number $\sigma(S)\geq\sigma,\;\forall S\in\mathcal{F}$in $\mathbb{R}^{d-1}$, then $\mathcal{F}$ contains a $(k-1)$-independent infinite sequence.
Inductive step:
Now we induct on $d$ and $k$. We have already proved the case for arbitrary $d\in\mathbb{N}$ and $k=0$, so we assume that $k>0$. If Claim 3.9 (1) is satisfied, then the theorem is proved in Lemma 3.11. So, let us assume that Claim 3.9 (2) is satisfied. Let $p$, $\vec{r}$ and $\mathcal{F}_{n}$ $\forall n\in\mathbb{N}$ be as defined in Claim 3.9 (2). Let $p$ be the origin $\mathcal{O}$ and $\vec{r}$ be the positive $d$-th axis. Then using Lemma 3.3, we can obtain a countably infinite family $\mathcal{G}\subseteq\mathcal{F}$ such that $\forall n\in\mathbb{N},\mathcal{G}\setminus\mathcal{F}_{n}$ is finite and $\mathcal{G}$ is a $k$-set. Therefore, $\mathcal{G}_{n}=\mathcal{G}\cap\mathcal{F}_{n}$ is a $k$-set, and by Claim 3.12, we can assume without loss of generality that for every $S\in\mathcal{G}_{n}$, $S\subset Q(\vec{r},n)$. Let $Y$ be the hyperplane perpendicular to $\vec{r}$ passing through $\mathcal{O}$. Then by the induction hypothesis, there exists an infinite sequence $\{S_{n}\}_{n\in\mathbb{N}}\subseteq\mathcal{G}$ whose projection on $Y$ forms a $(k-1)$-independent sequence. Since for every $n\in\mathbb{N}$, $\mathcal{G}\setminus\mathcal{F}_{n}$ is finite, we assume without loss of generality that $S_{n}\subset Q(\vec{r},n)$.
Now pick any distinct $k+1$ sets $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ from $\{S_{n}\}_{n\in\mathbb{N}}$. Let $\mathcal{C}_{k+1}(S_{i_{1}},\dots,S_{i_{k+1}})$ denote the union of all $k$-flats passing through $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$. Clearly, there is no $k$-flat that passes through $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ that contains a straight line parallel to $\vec{r}$. We claim that if every $k$-flat passing through $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ does not contain any straight line parallel to $\vec{r}$, then $\exists n^{\prime}\in\mathbb{N}$ such that $Q(\vec{r},n^{\prime})\cap\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})=\emptyset$. We shall prove this as Claim 3.13. Since there is a $S_{i_{k+2}}\subset Q(\vec{r},n^{\prime})$, we have $S_{i_{k+2}}\cap\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})=\emptyset$. In general, for $S_{i_{1}},\dots,S_{i_{m}}$ chosen from $\{S_{n}\}_{n\in\mathbb{N}}$ such that $\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})\cap S^{\prime}_{k+2}=\emptyset$ for any distinct $S^{\prime}_{1},\dots,S^{\prime}_{k+2}$ chosen from $S_{i_{1}},\dots,S_{i_{m}}$, by Claim 3.13 we can find an $n_{m}\in\mathbb{N}$ for which $Q(\vec{r},n_{m})\cap\mathcal{C}_{k+1}(S^{\prime}_{1},\dots,S^{\prime}_{k+1})=\emptyset$. We pick an $S_{i_{m+1}}$ from $\{S_{n}\}_{n\in\mathbb{N}}$ with $i_{m+1}>i_{j}$ for all $j\in[m]$ such that $S_{i_{m+1}}\subset Q(\vec{r},n_{m})$. Continuing this way, we obtain our required $k$-independent sequence. ∎
Now it remains to prove the following claim:
Claim 3.13.
If every $k$-flat passing through $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ does not contain any straight line parallel to $\vec{r}$, then $\exists n^{\prime}\in\mathbb{N}$ such that $Q(\vec{r},n^{\prime})\cap\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})=\emptyset$.
Proof.
Let, without loss of generality, $\vec{r}=\{ae_{d}\mid a\geq 0\}$, where $e_{d}=(0,\dots,0,1)$. Since no $k$-flat through $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ contains any straight line parallel to $\vec{r}$, we conclude that through any $c_{1},\dots,c_{k+1}$ with $c_{j}\in S_{i_{j}}$ for $j\in[k+1]$, exactly one $k$-flat passes.
Since the convex sets $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$ are compact and there is no $k$-flat parallel to $\vec{r}$ that intersects $S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}}$, we have $c>0$ such that $\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})\cap\{ae_{d}\in\mathbb{R}^{d}\;|\;a\geq c\}=\emptyset$. This we can show in the following way: define a function $\tau:S_{i_{1}}\times\dots\times S_{i_{k+1}}\to\mathbb{R}^{+}$ where $\tau(c_{1},\dots,c_{k+1})$ is the distance of the $k$-flat that passes through $c_{1},\dots,c_{k+1}$ from $\vec{r}$. As $\tau$ is continuous, $\tau^{-1}(\{0\})$ is closed and therefore compact in $S_{i_{1}}\times\dots\times S_{i_{k+1}}$. Let $C=\tau^{-1}(\{0\})$. Then if $\rho:C\to\mathbb{R}^{+}\cup\{0\}$ be the function for which $\rho(x_{1},\dots,x_{k+1})=a$ implies that the $k$-flat that passes through $x_{1},\dots,x_{k+1}$ intersects $\vec{r}$ at the point $ae_{d}$, then $\rho$ is continuous and therefore $\rho(C)$ is compact. From this, we get the required $c>0$.
Now we define $\vec{r^{\prime}}=ce_{d}+\vec{r}$ and $f:S_{i_{1}}\times S_{i_{2}}\times\dots\times S_{i_{k+1}}\rightarrow\mathbb{R}^{+}$ by the following: take any $c_{j}\in S_{j},\;\forall j\in[k+1]$ and suppose $\mathcal{K}$ be the $k$-flat passing through $c_{j}\;\forall j\in[k+1]$ and $u_{\mathcal{K}}\in\vec{r^{\prime}}$ be the point closest to $\mathcal{K}$. Then
$f(c_{i_{1}},\dots,c_{i_{k+1}})=\frac{\text{dist}(\mathcal{K},\vec{r^{\prime}})}{||u_{\mathcal{K}}||}.$
For any given $k$-flat $\mathcal{K}\in\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})$, $u_{\mathcal{K}}$ is uniquely defined as $\mathcal{K}$ contains no straight line parallel to $\vec{r}$. Also $u_{\mathcal{K}}\neq\mathcal{O}$, as $\mathcal{O}\not\in\vec{r^{\prime}}$. So $f$ is well-defined. Again $\forall(c_{i_{1}},\dots,c_{i_{k+1}})\in S_{i_{1}}\times S_{i_{2}}\times\dots\times S_{i_{k+1}}$, $f(c_{i_{1}},\dots,c_{i_{k+1}})\neq 0$. Observe that $f$ is continuous. So $\mathrm{image}(f)$ is compact. Hence $\exists n^{\prime}\in\mathbb{N}$ such that $f(S_{i_{1}}\times S_{i_{2}}\times\dots\times S_{i_{k+1}})\cap[0,\frac{1}{n^{\prime}}]=\emptyset$, and hence $\forall n>n^{\prime}(1+\left\lceil c\right\rceil)+1$ we have $Q(\vec{r},n)\cap\mathcal{C}_{k+1}(S_{i_{1}},S_{i_{2}},\dots,S_{i_{k+1}})=\emptyset$. ∎
4 Consequences of having strong colorful $(\aleph_{0},.)$-property
In this section, we shall prove Theorem 1.19 and Proposition 1.20.
Lemma 4.1.
For each $n\in\mathbb{N}$, suppose $\mathcal{F}_{n}$ is a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{n}$ is not pierceable by finite number of points. Then there exists a strictly heterochromatic sequence $\left\{B_{n}\right\}_{n\in\mathbb{N}}$ of $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ that is 00-independent.
Proof.
Our claim is that given any $R>0$ and any $n\in\mathbb{N}$, there exists $B\in\mathcal{F}_{n}$ such that $\text{dist}(B,O)>R$. Since there can be at most finitely many pairwise disjoint closed unit balls in $\mathcal{F}_{n}\cap B(O,R+1)$, we can choose $B\in\mathcal{F}_{n}$ with $\text{dist}(B,O)\geq R+1$.
Now suppose for any $m\in\mathbb{N}$, $B_{1},\dots,B_{m}$ are pairwise disjoint unit balls such that $\forall i\in[m],\;B_{i}\in\mathcal{F}_{i}$. Let $\bigcup_{i\in[m]}B_{i}\subseteq B(O,R_{m})$, for some $R_{m}>0$. Then by the above claim we can choose $B_{m+1}\in\mathcal{F}_{m+1}$ such that with $\text{dist}(B_{m+1},O)\geq R_{m+1}+1$. Then $\forall i\in[m],\;B_{i}\cap B_{m+1}=\emptyset$. Thus by induction on $m$, we can get the required heterochromatic sequence. ∎
See 1.19
Proof.
Using Lemma 4.1, we get that, there is at least one family $\mathcal{F}_{i}$, which is pierceable by a finite number of points. Now take any $B^{\prime}_{i}\in\mathcal{F}_{i}$ and for each $n\in\mathbb{N},\;n\neq i$, we define,
$\mathcal{F}^{\prime}_{n}=\{B\in\mathcal{F}_{n}\;|\;B\cap B^{\prime}_{i}=\emptyset\}.$
Then for each $n\in\mathbb{N},\;n\neq i,\;\mathcal{F}_{n}\setminus\mathcal{F}^{\prime}_{n}$ is pierceable by finite number of points. Then it is sufficient to show that at least one family $\mathcal{F}^{\prime}_{j},j\neq i$ is pierceable by a finite number of points. If $\mathcal{F}^{\prime}_{j}=\emptyset$, for some $j\neq i$, then we are done. Otherwise, $\forall j\in\mathbb{N},\;j\neq i,\;\mathcal{F}^{\prime}_{j}\neq\emptyset$. Now consider any heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}\setminus\{i\}}$ such that for each $n\in\mathbb{N}\setminus\{i\}$, $B_{n}\in\mathcal{F}^{\prime}_{n}$. Then $\{B_{n}\}_{n\in\mathbb{N}}$, where $B_{i}=B^{\prime}{{}_{i}}$ and $B_{n}\in\mathcal{F}^{\prime}_{n}$ for all $n\in\mathbb{N}\setminus\{i\}$, is a strict heterochromatic sequence and hence it contains at least one intersecting pair. Since for each $n\in\mathbb{N}\setminus\{i\},\;B_{n}\in\mathcal{F}^{\prime}_{n}$ implies $B_{n}\cap B^{\prime}_{i}=\emptyset$, so we must have at least one intersecting pair in $\{B_{n}\}_{n\in\mathbb{N}\setminus\{i\}}$. Using Lemma 4.1, we get that, there is at least one family $\mathcal{F}^{\prime}_{j}$, which is pierceable by a finite number of points. ∎
Figure 1: An illustration for the proof of Proposition 1.20.
Next, we shall show the tightness of the above result, that is Proposition 1.20. See 1.20
Proof.
Suppose $\mathcal{F}_{1}$ is the family containing only one unit ball $B_{1}$ in $\mathbb{R}^{2}$ with center at the origin and $\mathcal{F}_{2}$ contains the $4$ unit balls $B_{2,1},\dots,B_{2,4}$ in $\mathbb{R}^{2}$ with centers at the points $(1.5,0)$, $(-1.5,0)$, $(0,1.5)$, and $(0,-1.5)$ in $\mathbb{R}^{2}$ respectively. Then $\forall B\in\mathcal{F}_{2},\;B\cap B_{1}\neq\emptyset$ and clearly each of the families $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are pierceable by a finite number of points. Now for each $n\in\mathbb{N},\;n\geq 3$, we define $\mathcal{F}_{n}$ to be the family of unit balls $\{B_{n,j}\;|\;j\in\mathbb{N}\}$, where $B_{n,j}$ is the unit ball in $\mathbb{R}^{d}$ with center at $c_{n,j}=(4j,4n),\;\forall j\in\mathbb{N}$. Then for any $m,n,i,j\in\mathbb{N},\;m,n\geq 3$ with $(m,i)\neq(n,j)$ we have $B_{m,i}\cap B_{n,j}=\emptyset$ and for each $n\in\mathbb{N},\;n\geq 3$, the family $\mathcal{F}_{n}$ is not pierceable by finite number of points. Hence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ satisfies both the properties required. ∎
5 Impossibility results
In this section, we will prove Theorem 1.15, Theorem 1.16, Theorem 1.18 and Theorem 1.21.
See 1.15
Proof.
The construction here is similar to the construction in [KP22]. First we pick a ball $B_{1}=B(c_{1},r_{1})$. Next we pick another ball $F_{2}=B(c_{2},r_{2})$ such that $B_{1}\subset B(c_{2},\rho r_{2})$ and $B_{1}\cap F_{2}=\emptyset$. Let ¯$\bar{\ell}_{2}$ be a compact curve joining $c_{2}$ and any point of $B_{1}$. Define $B_{2}=F_{2}\cup\bar{\ell}_{2}$. In general, after $B_{1},B_{2},\dots,B_{m}$ are chosen, we pick another ball $F_{m+1}=B(c_{m+1},r_{m+1})$ such that $\cup_{i\in[m]}B_{i}\subset B(c_{m+1},\rho r_{m+1})$ and $\cup_{i\in[m]}B_{i}\cap F_{m+1}=\emptyset$. Let ¯$\bar{\ell}_{m+1}$ be a compact curve from $c_{m+1}$ that passes through $B_{i}$ for every $i\in[m]$ and does not intersect $B_{i}\cap B_{j}$ for any $i,j\in[m]$, $i\neq j$. Define $B_{m+1}=F_{m+1}\cup\bar{\ell}_{m+1}$ (see Figure 2). Then clearly, $\forall i\in[m],\;B_{m+1}\cap B_{i}\neq\emptyset$, i.e, $\{B_{n}\}_{n\in\mathbb{N}}$ satisfies the $(\aleph_{0},2)$ property. Again $\forall i,j\in[m]\text{ with }i\neq j,\;B_{m+1}\cap B_{i}\cap B_{j}=\emptyset$, i.e, no $3$ members of $\{B_{n}\}_{n\in\mathbb{N}}$ have a common point. So $\{B_{n}\}_{n\in\mathbb{N}}$ is not pierceable by a finite number of points. ∎
Figure 2: Sequence of non-convex sets with bounded condition number
See 1.16
We will first fix some notations that will be used in the proof of Theorem 1.16. Let’s fix some notations:
  • We will denote by $\mathbb{S}^{1}$ as the unit circle center at the origin in $\mathbb{R}^{2}$.
  • For all $n\in\mathbb{N}$, we will denote by $\theta_{n}$ the angle ${\pi}/{2^{n}}$.
  • Let $a=(0,1)$, $b=(1,0)$, $c=(1,1)$, and $u_{n}:=(\cos\theta_{n},\sin\theta_{n})$ for all $n\in\mathbb{N}$.
  • For all $n\in\mathbb{N}$, let $a_{n}$ and $b_{n}$ be the intersection points of the tangent to the circle $\mathbb{S}^{1}$ at the point $u_{n}$ with the lines $x_{2}=1$ and $x_{1}=1$ respectively.
  • Let :=$\mathcal{R}:=\left\{(\overline{x}_{1},\overline{x}_{2})\in\mathbb{R}^{2}\;\mid\;0\leq\overline{x}_{1}\leq 1\;\mbox{and}\;0\leq\overline{x}_{2}\leq 1\right\}$.
  • For all $n\in\mathbb{N}$, $s_{n}$ denotes the closed line segment joining the points $a_{n}$ and $b_{n}$. Let
    $B^{(i)}_{n}:=\left\{\overline{x}+\lambda u_{n}\;\mid\;\overline{x}\in s_{n},\,\lambda\in\left[0,1/2^{i}\right]\right\},$
    and $\mathcal{F}_{i}:=\left\{B_{n}^{(i)}\;\big{|}\;n\in\mathbb{N}\right\}$. See Figure 3.
Figure 3: The first figure shows the constructed line segments $s_{n}$, and the second figure shows the rectangles $B_{n}^{(i)}$.
Consider the family of rectangles
$\mathcal{F}:=\bigcup_{i\in\mathbb{N}}\mathcal{F}_{i}.$
Observe that for all $m$ and $n$ in $\mathbb{N}$, $s_{m}\cap s_{n}\neq\emptyset$, see Figure 4. Therefore, $B^{(i)}_{n}\cap B^{(j)}_{m}\neq\emptyset$. We will now show that $\mathcal{F}$ cannot be pierced by a finite set of points. To reach a contradiction suppose there exists a finite set $C\subsetneq\mathbb{R}^{2}$ such that for all $B_{n}^{(i)}\cap C\neq\emptyset$. Without loss of generality we can assume that the points $a=(0,1)$, $b=(1,0)$, and $c=(1,1)$ are in the set $C$.
First, we want to show that it will be sufficient to consider the case where
$C\subseteq\mathcal{R}\setminus B^{o}(O,1).$
Observe that by the construction of $\mathcal{F}$ if $C$ contains a point $p=(p_{1},p_{2})$ with $p_{1}<0$ or $p_{2}<0$ then $p$ does not hit any rectangles in $\mathcal{F}$. Therefore, we can assume that for all $p=(p_{1},p_{2})\in C$ we have $p_{1}\geq 0$ and $p_{2}\geq 0$. Now, if $C$ contains a point $p=(p_{1},p_{2})$ with $p_{1}>1$ then $C\cup\{\widetilde{p}\}\setminus\{p\}$ where $\widetilde{p}=(1,p_{2})$ is a point transversal of $\mathcal{F}$. Similarly, if $C$ contains a point $q=(q_{1},q_{2})$ with $q_{2}>1$ then $C\cup\{\widetilde{q}\}\setminus\{q\}$ where $\widetilde{q}=(q_{1},1)$ is also a point transversal of $\mathcal{F}$. Observe that none of the points in $B^{o}(O,1)$ can hit any rectangles in $\mathcal{F}$. Therefore, combining all the above observations we can safely assume that $C\subseteq\mathcal{R}\setminus B^{o}(O,1)$.
Figure 4: The first figure shows that any two segments always intersect, and the second figure shows how the rectangle $B_{n_{0}}^{(i_{0})}$ is disjoint from the set $C$.
Let $C_{1}$ be the subset of $C$ obtained by intersecting $C$ with the line $\ell_{1}:x_{1}-1=0$, and $C_{2}:=C\setminus C_{1}$. Also, let
$\delta:=\min_{q\in C_{2}}\text{dist}(q,\ell_{1}),\;\mbox{and}\;\lambda:=\min_{p\in C_{1}\setminus\{b\}}\|p-b\|.$
Observe that as
$\lim_{n\to\infty}a_{n}=(1,1)\;\mbox{and}\;\lim_{n\to\infty}b_{n}=(1,0),$
therefore there exists $n_{0}\in\mathbb{N}$ such that
$\|a_{n_{0}}-c\|<\delta\;\mbox{and}\;\|b_{n_{0}}-b\|<\lambda.$
See Figure 4.
Figure 5: The figure shows the triangles $T^{(i)}_{n}$ that can be used to establish Theorem 5.1.
Let $q_{1}$ be the point in the set $C_{1}\setminus\{b\}$ closest to $b$, and let $\ell_{n_{0}}$ be the line tangent to the circle $\mathbb{S}^{1}$ at the point $u_{n}$. Note that there exits $i_{0}\in\mathbb{N}$ such that
$1/2^{i_{0}}<\text{dist}(q_{1},\ell_{n_{0}}),$
see Figure 4. Now observe that the rectange $B_{n_{0}}^{(i_{0})}$ and the set $C$ is disjoint. Therefore, $C$ is not piercing $\mathcal{F}$ and we have reached a contradiction. This completes the proof of Theorem 1.16.
Observe that the construction can also be done along similar lines for non-axis parallel right-angled triangles in $\mathbb{R}^{2}$, and therefore we can prove the following result.
Theorem 5.1.
There exists a family $\mathcal{F}$ of pairwise intersecting right-angled triangles in $\mathbb{R}^{2}$ that cannot be pierced by any finite collection of points in $\mathbb{R}^{2}$. Note that the triangles in $\mathcal{F}$ have bounded diameters and non-zero volumes.
We will now prove Theorem 1.18.
Figure 6: One-unbounded family satisfying $(\aleph_{0},3)$-property with respect to lines without any finite line transversal.
See 1.18
Proof.
Recall the set of rectangles $\mathcal{F}$ from Theorem 1.16, where =$\mathcal{F}=\{R_{n}\}_{n\in\mathbb{N}}$. Now, for each $n\in\mathbb{N}$, lift $R_{n}$ to the plane $z=n$. Call the set of these lifted rectangles $\mathcal{F}^{\prime}$. Clearly, $\mathcal{F}^{\prime}$ does not have a finite hitting set consisting of lines. Now we show that every infinite subset of $\mathcal{F}^{\prime}$ has $3$ sets that are pierced by a single straight line. Let $\mathcal{S}\subset\mathcal{F}^{\prime}$ be an infinite subset of $\mathcal{F}^{\prime}$. For every $R\in\mathcal{S}$, choose a point $p_{R}\in R$. Let $q_{R}$ be the projection of $p_{R}$ onto the $xy$-plane. The set $\{q_{R}\}_{R\in\mathcal{S}}$ has a limit point $q$. Let $\vec{r}$ be a ray parallel to the $z$-axis with origin at $q$, where $\vec{r}$ is obtained by shifting the positive $z$-axis to $q$. For every $n\in\mathbb{N}$, $Q(\vec{r},n)$ (as defined in Section 2) contains $p_{R}$ for infinitely many $R\in\mathcal{S}$. Choose any two $R_{i},R_{j}\in\mathcal{S}$ and project them onto the $xy$-plane, and pick a point $q^{\prime}$ that is in the interior of the intersection of the two projected sets. Let $\vec{r}^{\prime}$ denote the ray from $q^{\prime}$ that is parallel to and is in the direction of $vec{r}$. Clearly, there is a cone $C$ with its endpoint at $q$ and $\vec{r}^{\prime}$ as its axis such that every ray in $C$ originating from $q^{\prime}$ intersects both $R_{i}$ and $R_{j}$ (see Figure 6). Since there is an $m\in\mathbb{N}$ for which $C\subset Q(\vec{r},m)$, the result follows. ∎
And we conclude this section by proving Theorem 1.21.
See 1.21
Figure 7: Dotted region denotes $A_{\{j_{1},j_{2}\}}$
Proof.
Suppose $\mathcal{I}^{\prime}$ is the collection of all multisets of size $(k+1)$ from $\mathbb{N}$, i.e, each $i\in\mathcal{I}^{\prime}$ is of the form $i=\{i_{1},\dots,i_{k+1}\}$ with each $i_{j}\in\mathbb{N}$ and all $i_{j}$’s are not necessarily distinct. Next suppose that $\mathcal{I}=[k+1]\cup\mathcal{I}^{\prime}$, i.e, $i\in\mathcal{I}$ implies either $i$ is a natural number lying in $[k+1]$ or $i$ is a multiset of size $(k+1)$ from $\mathbb{N}$. Then clearly $\mathcal{I}$ has a bijection with $\mathbb{N}$.
Now we take an infinite sequence $\{B_{n}\}_{n\in\mathbb{N}}$ of closed unit balls in $\mathbb{R}^{d}$ such that no $(k+2)$ balls from the sequence are pierceable by a single $k$-flat.
For each $\ell\in[k+1]$, define $\mathcal{F}_{\ell}$ to be the family $\{B_{n}\}_{n\in\mathbb{N}}$. And for any $i\in\mathcal{I}$ with the form $i=\{{j_{1}},\dots,j_{k+1}\}$, we define the region
$A_{i}:=\{x\in\mathbb{R}^{d}\;|\;\text{there is a $k$-flat passing through $x$ that pierces the balls }B_{j_{1}},\dots,B_{j_{k+1}}\},$
see Figure 7. Next we define the family $F_{i}$ to be the collection that we obtain by tightly packing the region $A_{i}$ with closed unit balls. Then $\mathcal{F}_{i}$ is not pierceable by finitely many $k$-flats. Thus for each $i\in\mathcal{I}$, we define $\mathcal{F}_{i}$ to be a family of closed unit balls in $\mathbb{R}^{d}$ such that $\mathcal{F}_{i}$ is not pierceable by a finite number of $k$-flats.
Now we choose any $k+1$ balls from the families $\mathcal{F}_{1},\dots,\mathcal{F}_{k+1}$, say, $B_{j_{1}},\dots,B_{j_{k+1}}$, where $B_{j_{\ell}}\in\mathcal{F}_{\ell},\forall\ell\in[k+1]$ and consider the $(k+1)$-tuple $i=\{{j_{1}},\dots,j_{k+1}\}$. Then $i\in\mathcal{I}$. Then by construction of the family $\mathcal{F}_{i}$, we get that each $B\in\mathcal{F}_{i}$ is pierceable by a $k$-flat that pierces all the $(k+1)$ balls $B_{j_{1}},\dots,B_{j_{k+1}}$. This proves the first part of the theorem. ∎